### v1

# AGN-heated dust revealed in “Little Red Dots”

AGN-heated dust revealed in “Little Red Dots”
Little Red Dots (LRDs) are a puzzling population of extragalactic sources whose origin is highly debated. In this Letter, we stack NIRCam, MIRI and ALMA images of a large and homogeneously-selected sample of LRDs from multiple JWST Legacy fields. We find clear evidence for hot-dust emission, seen as a rising mid-infrared continuum up to rest-frame 3m, that is best explained by a standard dusty AGN structure. This scenario holds regardless of whether the Optical/Ultraviolet (UV) continuum is stellar or AGN-dominated. Either ways, we argue that AGN-heated dust must have a geometry that leaves at least part of the Optical/UV continuum and broad-line region emission unobscured, which explains the observed blue UV slopes and prominent Balmer features. Lack of detections in the deep stack of X-ray images suggests that Compton-thick (31024 cm-2) gas obscuration is common, and likely confined within the dust sublimation radius (sub0.1 pc). While this gas-dust displacement is in line with recent models (e.g. the “Black Hole Star”), how pre-enriched hot dust can form around a nearly-pristine black hole environment remains to be explained.
Key Words.:
galaxies: nuclei – galaxies: active – galaxies: evolution1 Introduction
The James Webb Space Telescope (JWST) has ushered extragalactic astronomy in a transformative era, offering unprecedented access to the high-redshift universe. One of the most puzzling discoveries is a population of sources dubbed as “Little Red Dots” (LRDs, e.g. Matthee et al. 2024, Greene et al. 2024, Labbe et al. 2025). LRDs are extremely compact (often unresolved at 100 pc scales) and show red optical colors combined with blue UV slopes, resulting in a characteristic “V-shaped” spectral energy distribution (SED), with a sharp turnover near the Balmer break at rest Å (e.g. Setton et al. 2024, Hainline et al. 2025). Remarkably, LRDs appear to be two orders of magnitude more abundant than luminous quasars at similar redshifts (Kokorev et al. 2024, Akins et al. 2024), making them accessible even in narrow-field JWST surveys.
However, an observational consensus of the physical nature of LRDs is still lacking. The red optical colors and broad emission lines point toward obscured Active Galactic Nuclei (AGN) activity (Inayoshi et al. 2020, Volonteri et al. 2021, Ji et al. 2025). In this case, the AGN continuum would be moderately reddened by dust (– mag; Casey et al. 2024; Akins et al. 2024; Wang et al. 2025). Yet, this interpretation is apparently at odds with their extreme X-ray and radio weakness implied by the lack of stacked X-ray (Ananna et al. 2024; Yue et al. 2024; Maiolino et al. 2025) and radio emission (Akins et al. 2024; Gloudemans et al. 2025). To alleviate this tension, some theoretical models postulate LRDs as supermassive black holes (SMBHs) accreting at super-Eddington luminosities and enshrouded in a turbulent, dust-free, and highly dense atmosphere: the “BH-star” scenario. This explains the strong Balmer break and the red continuum through pure gas opacity, and the broad Balmer lines via resonant electron scattering (e.g. Naidu et al. 2025; Inayoshi & Maiolino 2025; Rusakov et al. 2025; de Graaff et al. 2025). Similar gas conditions are postulated in other models, with the presence of a metal-free halo surrounding a pre-collapse “super-massive star” (Nandal & Loeb 2025), or a late-stage “Quasi star” (Begelman & Dexter 2025). All these models share the absence of prior metal enrichment and the presence of a gas-dense, dust-free region at sub-pc scales. So far, this dust paucity is supported by the flat LRD SED observed beyond rest 1.6 m (e.g. Williams et al. 2024; Carranza-Escudero et al. 2025), in contrast to the rising mid-infrared (MIR) continuum expected in obscured AGN, conveying the idea that MIR emission in LRDs is dominated by dust-poor stellar populations and not AGN dust (Williams et al. 2024; Pérez-González et al. 2024, Leung et al. 2024). In addition, the non-detection from deep ALMA stacking poses a stringent dust mass limit of Mdust106 M⊙ at 6 (Casey et al. 2025; see also Chen et al. 2025; Setton et al. 2025). Nevertheless, we caution that previous NIRCam/MIRI SED studies of LRDs might have been biased towards bright objects (e.g. Labbe et al. 2024; Barro et al. 2024; Chen et al. 2025; de Graaff et al. 2025; Wang et al. 2025), or possibly affected either by individual MIRI non-detections (e.g. Pérez-González et al. 2024) or by stacking peculiar LRD sub-samples (e.g. -dropouts; Williams et al. 2024), whose SED could intrinsically differ from that of “typical” LRDs. To explore this possibility, in this Letter we stack a large and homogeneously-selected sample of LRDs with broad NIRCam/MIRI coverage. Throughout, we assume a Chabrier (2003) IMF and a flat CDM cosmology with = 70 km/s/Mpc, , and .
2 Photometric selection of LRDs
We select LRDs from the latest compilation of Kocevski et al. (2025), consisting of 341 photometrically-selected LRDs in the redshift range 2–11, and retrieved from several JWST/NIRCam surveys. These LRDs were homogeneously selected to display compact morphologies and “V-shaped” continuum from spectral slope fitting between blueward and redward of the rest-frame 3645 Å (i.e. close to the Balmer break). These criteria enable a self-consistent search for LRD candidates with red optical and blue UV colors over a wide redshift interval (Hainline et al. 2025), although we acknowledge that it does not guarantee an homogeneous physical selection of LRDs.
We replace photometric redshifts () with NIRSpec-based spectroscopic redshifts () taken from the public Dawn JWST Archive (DJA; Brammer 2023b; de Graaff et al. 2024; Heintz et al. 2024), whenever of high-quality (grade=3, see archive). This increases the spectroscopic fraction from 11% (39/341) to 29% (100/341). The mutual agreement between and estimates is generally good, with only five sources showing an offset /(1+)0.2.
Finally, we restrict ourselves to the NIRCam fields with (at least partial) MIRI coverage in two or more bands (including 18m and/or 21m; see Table 1), which yield the strongest IR-SED constraints. Our final sample counts 302 LRDs taken from the following surveys: the Cosmic Evolution Early Release Science Survey (CEERS; Finkelstein et al. 2023) in the EGS; the JWST Advanced Deep Extragalactic Survey (JADES; Eisenstein et al. 2023) in the GOODS-S, complemented by MIRI imaging from the Systematic Mid-infrared Instrument Legacy Extragalactic Survey (SMILES; Rieke et al. 2024; Alberts et al. 2024); the Public Release IMaging for Extragalactic Research (PRIMER111https://primer-jwst.github.io/) survey in the COSMOS and UDS fields. Out of 302 LRDs, 163 (54%) are covered by MIRI 7.7m, 126 (42%) also by MIRI-18m, and 85 (28%) have a robust . Moreover, we exploit archival ALMA/B6 (1.1 mm) data from the combined A3COSMOS+A3GOODSS surveys (Liu et al. 2019; Adscheid et al. 2024; Magnelli et al. 2024), which are available for a total of 70 LRDs in the COSMOS and GOODS-S fields. The number of sources split down by field, photometric coverage and spectroscopic redshifts is summarized in Table 1.
3 Stacking results
We perform median stacking over all available NIRCam bands (7) and all MIRI bands (2 in PRIMER, 7 in CEERS and 8 in SMILES), whenever available, at the optical position of all LRDs. We use a fixed pixel scale of 0.06′′, and we extract fluxes from a circular aperture with fixed radius =0.25′′. ALMA mean stacking, instead, is performed in the -plane, in order to homogeneously combine archival data at different frequency, angular resolution and sensitivity. Further details of the stacking methodology are given in Appendix A. We refer the reader to Appendix B, to verify the robustness of our stacking results against possible biases. In Fig. 1 we show the stacked cutouts (3′′3′′) in each band, sorted by increasing wavelength. For comparison, the dashed white circle marks the corresponding PSF FWHM. As MIRI images are strongly oversampled, to aid visualization we smooth them with a Gaussian kernel, adopting a larger radius for larger PSF FWHM (see caption of Fig. 1). The stacked ALMA/B6 image is obtained in the -plane and imaged at 0.1′′/px scale. Except for MIRI/f2550w and ALMA/B6, all other median stacks lead to a detection with signal-to-noise ratio (S/N) greater than 3. We stress that all detections are consistent with being unresolved, as their extent is smaller or comparable with the PSF FWHM of the corresponding band (white circles in Fig. 1). All fluxes, errors, rms of the stacked images are listed in Table 2 along with the effective number of sources stacked in each band.
4 Discussion and Conclusions
4.1 SED-fitting with vs without AGN
In order to interpret the stacked photometry of LRDs, we perform SED-fitting using the Code Investigating GALaxy Emission (CIGALE; Boquien et al. 2019). We fix the input redshift to the median value of the LRD sample (6.2). In order to account for the broad redshift distribution of the sample (mostly at 4.58; see Fig. 4) and the correspondingly wide rest-frame range sampled by each filter, we convolve the filter response curves at 6.2 with the band-specific redshift distribution of the underlying LRD sample. Then, we run CIGALE both with and without AGN templates to test the relative significance of the AGN component. In both SED runs, the library of stellar templates is taken from Bruzual & Charlot (2003), including a delayed star formation history with random bursts. Energy balance is assumed between the dust-absorbed UV radiation and the reprocessed IR emission, modeled via a dust attenuation curve from Charlot & Fall (2000), and dust templates from Dale et al. (2014). The AGN templates come from the skirtor module (Stalevski et al. 2012; 2016), which includes emission from an accretion disk, a dusty torus and polar dust. A comprehensive list of the SED-fitting parameters is given in Appendix C and D. In Fig. 2, we display all stacked fluxes (or 3 upper limits) overlaid to the total best-fit SED (black line) obtained from the run without AGN (top panel) or with AGN (middle panel).
We caution the reader that our SED-fitting analysis is not primarily aimed at interpreting the rest optical or far-IR properties of LRDs. For instance, CIGALE assumes a single-zone obscuration between optical, UV and IR light, which might be oversimplistic in case of no co-spatiality (see Sect. 4.2). Nonetheless, regardless of the origin of Optical/UV emission in LRDs, we argue that AGN-heated dust is needed to reproduce the monothonic increase up to rest-frame 2–3m, which pure dust-attenuated stellar light fails to match. This is further confirmed by our ALMA upper limit, which is very similar to that derived by Casey et al. (2025) (see also Xiao et al. 2025; Setton et al. 2025), which translates into Mdust106 M⊙ for T100 K. Indeed, the best-fit SED without AGN strongly overshoots the ALMA upper limit when attempting to reproduce the power-law-like SED at rest 1–3m. While this could be resolved by lifting the energy balance implemented in CIGALE, even ignoring the upper limits (at f2550w and ALMA/B6), a simple F-test comparing the two reduced values favours the AGN solution at 95% significance.
4.2 Are LRDs typically dust-obscured AGN?
By considering the best-fit solution with AGN, the rest-frame 6 m luminosity of the AGN torus can be used as a proxy for the intrinsic AGN X-ray (rest [2–10] keV) luminosity (X,intr (e.g. Stern 2015). From 6μm=(3.80.6)1043 erg/s, we derive X,intr=21043 erg/s, which nicely matches that expected from the best-fit AGN bolometric luminosity, if assuming standard X-ray bolometric corrections (e.g. Lusso et al. 2012). By stacking the deepest X-ray images of the 44 LRDs falling inside the Chandra Deep Field South (Luo et al. 2017), we obtain a 3 [2–10] keV luminosity limit X1042 erg/s at =6.2 (Comastri et al. subm). In order to match it with the X,intr expected from the AGN SED, the hydrogen column density has to be NH31024 cm-2, implying Compton-thick gas obscuration (Ananna et al. 2024; Yue et al. 2024; Maiolino et al. 2025).
Given the SED-based estimate of the AGN bolometric luminosity, erg s-1 and the best-fit stellar mass from the AGN solution (M⋆2.5109 M⊙; i.e. 2 smaller than from a pure-galaxy fit), we emphasize that, if LRDs are highly-accreting AGN (Eddington ratio 10%), the BH-to-galaxy mass ratio MBH/M⋆1:100, implying strongly overmassive BHs compared to local scaling relations (see also Pacucci et al. 2023; Juodžbalis et al. 2024; Maiolino et al. 2024). Moreover, our estimate corresponds to a dust sublimation radius of 0.1 pc at K (Barvainis 1987), meaning that any dusty AGN structure should begin at scales pc from the central SMBH. In the bottom panel of Fig. 2, we try to decompose the AGN-dust template with three multi-temperature greybody curves (with free to vary). Our rising MIRI SED up to rest-frame 2–3m indicates hot dust at K, though it could be even hotter if optically-thick. Instead, the bulk dust mass (Mdust3104 M⊙) comes from colder polar AGN dust at 50 K (though not constrained by our data). While a dusty structure could induce X-ray obscuration, it would also strongly suppress the Optical/UV continuum, in tension with the blue UV slope and strong Balmer features that imply relatively modest dust attenuation (; Akins et al. 2024; Casey et al. 2025).
This apparent contradiction can be resolved if X-ray absorption in LRDs occurs within , in a gas-dense and dust-free region. This gas-dust displacement is formally consistent with the recently-proposed models of Quasi-star (Begelman & Dexter 2025), BH–star (Naidu et al. 2025), and super-massive star (Nandal & Loeb 2025) scenarios. However, the presence of pre-enriched dust around a first-generation (or metal-poor) star is hard to justify. Moreover, the normalized BH–star SED (green line in Fig. 2, bottom panel) appears to flatten out at rest-frame 1 m, contrary to the rising MIRI fluxes.
Our findings might fit better with an alternative hypothesis suggested by previous studies (Setton et al. 2024; Maiolino et al. 2025; D’Eugenio et al. 2025; Ji et al. 2025; Inayoshi & Maiolino 2025), in which X-ray obscuration arises from Compton-thick gas clouds within the BLR (at sub, i.e. dust-free), which covers the compact X-ray corona without suppressing the broad Balmer lines. Yet, the origin of the Optical/UV continuum remains open. Although we caution that LRDs might not consist of a single population, we speculate two simple scenarios in Fig. 3, depending on what powers the Optical/UV continuum in LRDs. In either cases, the bulk of Optical/UV continuum is not dust-attenuated, despite the presence of hot AGN dust, and the outer BLR is directly visible along the line of sight. (i) If the Optical/UV comes from stellar light (Fig. 3, left), the LRD resembles a “type-2” AGN configuration, in which the torus covers the AGN accretion disk, which would otherwise outshine the host. In case of high dust clumpiness, mildly-attenuated light from the outer accretion disk could add to the optical emission. (ii) Instead, if the Optical/UV comes from AGN light (Fig. 3, right), the LRD resembles a “type-1” unobscured AGN configuration, though the origin of the sharp Balmer break remains unclear. Fainter Optical/UV light from kpc-scale star formation cannot be ruled out (e.g. Rinaldi et al. 2025).
In summary, we find clear evidence for a rising MIRI continuum up to 3m, which we interpret via SED fitting as hot (8205 K) AGN dust, regardless of whether the Optical/UV continuum is stellar- or AGN-dominated, or a combination thereof. Despite the presence of hot dust, these LRDs selected via V-shaped and compactness criteria must be not strongly dust-obscured. We acknowledge that such homogeneous selection does not necessarily translate into a single LRD population with common physical properties. Our findings mainly provide en empirical benchmark for the presence of AGN-heated dust in the majority of LRDs, or (at least) in a significant sub-population of them. For instance, if LRDs are standard AGN, we speculate that the above selection criteria might be biased against the most dust-obscured (“type-2”) LRD analogs, which could be recovered by lifting the blue-UV slope requirement.
Acknowledgements.
ID acknowledges funding by the European Union – NextGenerationEU, RRF M4C2 1.1, Project 2022JZJBHM: ”AGN-sCAN: zooming-in on the AGN-galaxy connection since the cosmic noon” - CUP C53D23001120006. (Some of) the data producs presented here were retrieved from the DJA, an initiative of the Cosmic Dawn Center (DAWN), which is funded by the Danish National Research Foundation under grant DNRF140.References
- Adscheid et al. (2024) Adscheid, S., Magnelli, B., Liu, D., et al. 2024, A&A, 685, A1
- Akins et al. (2024) Akins, H. B., Casey, C. M., Lambrides, E., et al. 2024, arXiv e-prints, arXiv:2406.10341
- Alberts et al. (2024) Alberts, S., Lyu, J., Shivaei, I., et al. 2024, ApJ, 976, 224
- Ananna et al. (2024) Ananna, T. T., Bogdán, Á., Kovács, O. E., Natarajan, P., & Hickox, R. C. 2024, ApJ, 969, L18
- Barro et al. (2024) Barro, G., Perez-Gonzalez, P. G., Kocevski, D. D., et al. 2024, arXiv e-prints, arXiv:2412.01887
- Barvainis (1987) Barvainis, R. 1987, ApJ, 320, 537
- Begelman & Dexter (2025) Begelman, M. C. & Dexter, J. 2025, arXiv e-prints, arXiv:2507.09085
- Boquien et al. (2019) Boquien, M., Burgarella, D., Roehlly, Y., et al. 2019, A&A, 622, A103
- Brammer (2023a) Brammer, G. 2023a
- Brammer (2023b) Brammer, G. 2023b
- Bruzual & Charlot (2003) Bruzual, G. & Charlot, S. 2003, MNRAS, 344, 1000
- Carranza-Escudero et al. (2025) Carranza-Escudero, M., Conselice, C. J., Adams, N., et al. 2025, arXiv e-prints, arXiv:2506.04004
- Casey et al. (2025) Casey, C. M., Akins, H. B., Finkelstein, S. L., et al. 2025, arXiv e-prints, arXiv:2505.18873
- Casey et al. (2024) Casey, C. M., Akins, H. B., Kokorev, V., et al. 2024, ApJ, 975, L4
- Chabrier (2003) Chabrier, G. 2003, PASP, 115, 763
- Charlot & Fall (2000) Charlot, S. & Fall, S. M. 2000, ApJ, 539, 718
- Chen et al. (2025) Chen, K., Li, Z., Inayoshi, K., & Ho, L. C. 2025, arXiv e-prints, arXiv:2505.22600
- Dale et al. (2014) Dale, D. A., Helou, G., Magdis, G. E., et al. 2014, ApJ, 784, 83
- de Graaff et al. (2024) de Graaff, A., Rix, H.-W., Carniani, S., et al. 2024, A&A, 684, A87
- de Graaff et al. (2025) de Graaff, A., Rix, H.-W., Naidu, R. P., et al. 2025, arXiv e-prints, arXiv:2503.16600
- D’Eugenio et al. (2025) D’Eugenio, F., Maiolino, R., Perna, M., et al. 2025, arXiv e-prints, arXiv:2503.11752
- Eisenstein et al. (2023) Eisenstein, D. J., Johnson, B. D., Robertson, B., et al. 2023, arXiv e-prints, arXiv:2310.12340
- Finkelstein et al. (2023) Finkelstein, S. L., Bagley, M. B., Ferguson, H. C., et al. 2023, ApJ, 946, L13
- Franco et al. (2025) Franco, M., Casey, C. M., Koekemoer, A. M., et al. 2025, arXiv e-prints, arXiv:2506.03256
- Gloudemans et al. (2025) Gloudemans, A. J., Duncan, K. J., Eilers, A.-C., et al. 2025, ApJ, 986, 130
- Greene et al. (2024) Greene, J. E., Labbe, I., Goulding, A. D., et al. 2024, ApJ, 964, 39
- Hainline et al. (2025) Hainline, K. N., Maiolino, R., Juodžbalis, I., et al. 2025, ApJ, 979, 138
- Harish et al. (2025) Harish, S., Kartaltepe, J. S., Liu, D., et al. 2025, arXiv e-prints, arXiv:2506.03306
- Heintz et al. (2024) Heintz, K. E., Watson, D., Brammer, G., et al. 2024, Science, 384, 890
- Inayoshi & Maiolino (2025) Inayoshi, K. & Maiolino, R. 2025, ApJ, 980, L27
- Inayoshi et al. (2020) Inayoshi, K., Visbal, E., & Haiman, Z. 2020, ARA&A, 58, 27
- Ji et al. (2025) Ji, X., Maiolino, R., Übler, H., et al. 2025, arXiv e-prints, arXiv:2501.13082
- Juodžbalis et al. (2024) Juodžbalis, I., Maiolino, R., Baker, W. M., et al. 2024, Nature, 636, 594
- Kocevski et al. (2025) Kocevski, D. D., Finkelstein, S. L., Barro, G., et al. 2025, ApJ, 986, 126
- Kokorev et al. (2024) Kokorev, V., Caputi, K. I., Greene, J. E., et al. 2024, ApJ, 968, 38
- Labbe et al. (2025) Labbe, I., Greene, J. E., Bezanson, R., et al. 2025, ApJ, 978, 92
- Labbe et al. (2024) Labbe, I., Greene, J. E., Matthee, J., et al. 2024, arXiv e-prints, arXiv:2412.04557
- Leung et al. (2024) Leung, G. C. K., Finkelstein, S. L., Pérez-González, P. G., et al. 2024, arXiv e-prints, arXiv:2411.12005
- Liu et al. (2019) Liu, D., Schinnerer, E., Groves, B., et al. 2019, ApJ, 887, 235
- Luo et al. (2017) Luo, B., Brandt, W. N., Xue, Y. Q., et al. 2017, ApJS, 228, 2
- Lusso et al. (2012) Lusso, E., Comastri, A., Simmons, B. D., et al. 2012, MNRAS, 425, 623
- Magnelli et al. (2024) Magnelli, B., Adscheid, S., Wang, T.-M., et al. 2024, A&A, 688, A55
- Maiolino et al. (2025) Maiolino, R., Risaliti, G., Signorini, M., et al. 2025, MNRAS, 538, 1921
- Maiolino et al. (2024) Maiolino, R., Scholtz, J., Curtis-Lake, E., et al. 2024, A&A, 691, A145
- Matthee et al. (2024) Matthee, J., Naidu, R. P., Brammer, G., et al. 2024, ApJ, 963, 129
- Naidu et al. (2025) Naidu, R. P., Matthee, J., Katz, H., et al. 2025, arXiv e-prints, arXiv:2503.16596
- Nandal & Loeb (2025) Nandal, D. & Loeb, A. 2025, arXiv e-prints, arXiv:2507.12618
- Pacucci et al. (2023) Pacucci, F., Nguyen, B., Carniani, S., Maiolino, R., & Fan, X. 2023, ApJ, 957, L3
- Pérez-González et al. (2024) Pérez-González, P. G., Barro, G., Rieke, G. H., et al. 2024, ApJ, 968, 4
- Rieke et al. (2024) Rieke, G. H., Alberts, S., Shivaei, I., et al. 2024, ApJ, 975, 83
- Rinaldi et al. (2025) Rinaldi, P., Rieke, G. H., Wu, Z., et al. 2025, arXiv e-prints, arXiv:2507.17738
- Rusakov et al. (2025) Rusakov, V., Watson, D., Nikopoulos, G. P., et al. 2025, arXiv e-prints, arXiv:2503.16595
- Setton et al. (2024) Setton, D. J., Greene, J. E., de Graaff, A., et al. 2024, arXiv e-prints, arXiv:2411.03424
- Setton et al. (2025) Setton, D. J., Greene, J. E., Spilker, J. S., et al. 2025, arXiv e-prints, arXiv:2503.02059
- Stalevski et al. (2012) Stalevski, M., Fritz, J., Baes, M., Nakos, T., & Popović, L. Č. 2012, MNRAS, 420, 2756
- Stalevski et al. (2016) Stalevski, M., Ricci, C., Ueda, Y., et al. 2016, MNRAS, 458, 2288
- Stern (2015) Stern, D. 2015, ApJ, 807, 129
- Taylor et al. (2025) Taylor, A. J., Kokorev, V., Kocevski, D. D., et al. 2025, ApJ, 989, L7
- Valentino et al. (2023) Valentino, F., Brammer, G., Gould, K. M. L., et al. 2023, ApJ, 947, 20
- Volonteri et al. (2021) Volonteri, M., Habouzit, M., & Colpi, M. 2021, Nature Reviews Physics, 3, 732
- Wang et al. (2025) Wang, B., de Graaff, A., Davies, R. L., et al. 2025, ApJ, 984, 121
- Wang et al. (2024) Wang, T.-M., Magnelli, B., Schinnerer, E., et al. 2024, A&A, 681, A110
- Wang et al. (2022) Wang, T.-M., Magnelli, B., Schinnerer, E., et al. 2022, A&A, 660, A142
- Williams et al. (2024) Williams, C. C., Alberts, S., Ji, Z., et al. 2024, ApJ, 968, 34
- Xiao et al. (2025) Xiao, M., Oesch, P. A., Bing, L., et al. 2025, arXiv e-prints, arXiv:2503.01945
- Yue et al. (2024) Yue, M., Eilers, A.-C., Ananna, T. T., et al. 2024, ApJ, 974, L26
Appendix A Stacked samples and fluxes
Here we provide a more detailed outline of the input images and stacking technique used in this work.
Image and noise maps are retrieved from the corresponding mosaics of each field (see below). In this step, we discard narrow or medium-band filters (whenever available) to mitigate potential flux boosting in the case of strong emission lines.
We retrieve PRIMER-COSMOS images as part of the COSMOS-Web team, where the full PRIMER-COSMOS mosaic has been reduced following the same procedure as with the COSMOS-Web data (Franco et al. 2025; Harish et al. 2025). All NIRCam and MIRI mosaics have a pixel size of 0.06′′.
For the GOODS-S, we retrieve NIRCam and MIRI images from the JADES and SMILES surveys publicly available on the MAST222SMILES: https://archive.stsci.edu/hlsp/smiles; JADES: https://archive.stsci.edu/hlsp/jades archive. While SMILES maps are already at 0.06′′/px, JADES maps (with a native pixel scale of 0.03′′) are re-scaled to 0.06′′/px for consistency with the other fields.
For CEERS and PRIMER-UDS, which are unavailable on MAST, we leverage the DJA (Brammer 2023a; Valentino et al. 2023), where the full reduced mosaics are publicly available in each NIRCam and MIRI band333DJA full reduced mosaics in each band are taken from https://s3.amazonaws.com/grizli-v2/JwstMosaics/v7/index.html, all with a pixel scale of 0.04′′ and in units of [10 nJy]. We do a resampling of all mosaics to 0.06′′/px for consistency with SMILES and PRIMER-COSMOS maps, while preserving astrometry and total flux over a given aperture. For PRIMER-UDS, we also merge the East and West mosaics (listed separately in the DJA) into a full mosaic.
We only consider sources whose position in the map is at least 20 pixels inside the edge of the mosaic. For each object and in each band, we create a pixel image (with =51, 0.06′′/px) from the corresponding map, each centred on the NIRCam position of the LRD. We then stack all stack sources and retrieve the median flux at each pixel.
The uncertainty on the stacked flux density is quantified through a bootstrapping technique: being stack the number of input targets in a given band, in each random realisation we re-shuffle the input sample, preserving the same stack by allowing source duplication. We repeat this 100 times, and take the median of the resulting flux distribution as our formal stacked flux. The 1 dispersion around this value is interpreted as the flux error. The rms of the stacked image is computed by stacking the inverse-variance weighted noise map of each target. Our median stacks yield a detection in each band up to 21 m with signal-to-noise444We consider as “noise” the nominal rms of the stacked image; though using the flux error derived via bootstrapping does not change our conclusions. S/N3 (see Table 2).
For ALMA, we perform a stacking analysis in the uv-domain using all publicly available Band-6 observations covering our LRDs in the COSMOS and GOODS-S fields. This analysis is carried out with CASA, following the method described in Wang et al. (2022) (see also Wang et al. 2024; Magnelli et al. 2024), to which we refer for further details of the procedure.
Fig. 4 displays the redshift distribution of our final LRD sample (302 sources, grey dashed), split by field (coloured open histograms). Filled histograms highlight the spectroscopic subsets. Corresponding numbers are listed in Table 1.
| Survey/Field | Instrument | Filters | Area | stack range |
| (arcmin2) | (spec) | |||
| CEERS/EGS | NIRCam | [f090w, f115w, f150w, f200w, f277w, f356w, f444w] | 88.6 | 63 (30) |
| MIRI | [f560w, f770w, f1000w, f1280w, | |||
| f1500w, f1800w, f2100w] | 55.0 | 2–39 (2–19) | ||
| JADES/GOODSS | NIRCam | [f090w, f115w, f150w, f200w, f277w, f356w, f444w] | 62.1 | 44 (7) |
| SMILES/GOODSS | MIRI | [f560w, f770w, f1000w, f1280w, | ||
| f1500w, f1800w, f2100w, f2550w] | 34.0 | 22–23 (2) | ||
| A3-GOODSS | ALMA | Band-6 (1080m) | 86.8 | 34 (4) |
| PRIMER/COSMOS | NIRCam | [f090w, f115w, f150w, f200w, f277w, f356w, f444w] | 138.9 | 79–80 (7) |
| MIRI | [f770w, f1800w] | 73.0 | 41–42 (3) | |
| A3-COSMOS | ALMA | Band-6 (1080m) | 127.9 | 36 (2) |
| PRIMER/UDS | NIRCam | [f090w, f115w, f150w, f200w, f277w, f356w, f444w] | 243.0 | 116 (41) |
| MIRI | [f770w, f1800w] | 130.0 | 61–62 (23) |
| wavelength | stack | rms | Flux | Error Flux |
| (m) | (mJy) | (mJy) | (mJy) | |
| 0.90 | 302 | 3.47e-09 | 8.49e-06 | 9.31e-07 |
| 1.15 | 302 | 4.64e-09 | 1.31e-05 | 1.06e-06 |
| 1.50 | 302 | 8.55e-09 | 1.47e-05 | 1.07e-06 |
| 2.00 | 302 | 1.32e-08 | 1.78e-05 | 1.39e-06 |
| 2.77 | 302 | 1.10e-08 | 3.03e-05 | 2.13e-06 |
| 3.56 | 302 | 1.87e-08 | 5.89e-05 | 4.41e-06 |
| 4.44 | 302 | 3.84e-08 | 1.13e-04 | 7.17e-06 |
| 5.60 | 28 | 3.20e-06 | 1.23e-04 | 3.06e-05 |
| 7.70 | 163 | 2.07e-06 | 2.13e-04 | 1.60e-05 |
| 10.00 | 61 | 1.07e-05 | 1.94e-04 | 3.35e-05 |
| 12.80 | 24 | 3.93e-05 | 2.68e-04 | 5.67e-05 |
| 15.00 | 61 | 3.35e-05 | 3.42e-04 | 5.73e-05 |
| 18.00 | 126 | 4.99e-05 | 5.33e-04 | 9.97e-05 |
| 21.00 | 60 | 1.30e-04 | 6.08e-04 | 1.36e-04 |
| 25.50 | 23 | 1.69e-03 | ¡5.07e-03 | — |
| 1080.00 | 70 | 5.89e-03 | ¡1.77e-02 | — |
Appendix B Stacking tests
We test our stacking results against multiple potential biases related to sample selection, stacking method, field-to-field variations and photometric redshift uncertainties. The main outcome of these tests is shown in Fig. 5. In general, these tests further corroborate the robustness of our stacking results and the need for an obscured AGN component to reproduce the MIR excess. Each of the following tests is discussed in a dedicated subsection, numbered as the panels of Fig. 5 from top to bottom.
B.1 Mean vs median stacking
In the first panel of Fig. 5 we compare our median stacked fluxes (grey open circles) against mean stacked fluxes (red starred symbols). These latter values are computed from the inverse-variance weighted flux density map, accounting for the different sensitivity (i.e. exposure time) of each target. Mean stacking leads to higher fluxes with respect to median stacking. This is likely caused by the fact that the resulting (linear) mean flux can be boosted upwards by relatively bright individual detections (we note that upper limits at f2550w and ALMA/B6 are unchanged as they are calculated from the rms map). This is the case in NIRCam bands, where all LRDs are individually detected by definition over multiple bands, both blueward and redward the Balmer break (in order to assess the V-shape). A similar effect is seen at MIRI 7.7m and 18m, which include about 100 more targets than at other MIRI bands (see Table 1), thanks to a complete coverage across all the fields studied here. Even a handful of bright detections at 7.7m and 18m (some have S/N100 in the PRIMER fields) are able to boost the mean MIRI flux upwards. Moreover, we opt for median stacking also because it strongly mitigates contamination from bright neighbors and catastrophic outliers, and thus reduces the confusion noise for the faint sources. This comparison, however, demonstrates that the mean stacked LRD SED reveals a similar trend as the median stacked SED, showing a rising MIR slope that is indicative of dusty AGN activity.
B.2 Field-to-field variations
An important aspect that our analysis relies on is the combination of multiple NIRCam and MIRI survey fields. While this is necessary to improve the statistics and the stacked S/N, it might potentially introduce systematics related to different JWST coverage and depth, which could affect the properties of photometrically-selected LRDs. We thus verify that our stacking results are robust against field-to-field variations, by repeating the same analysis separately for each field. The second panel in Fig. 5 displays the median stacked SED obtained in each field (represented by different filled symbols) overlaid to our full stacking results (grey open circles). For visual purposes, here we relax the detection threshold to S/N2 (hence upper limit fluxes are set to twice the noise level if S/N2). Despite their heterogeneous spectral coverage across MIRI bands, all single-field median SEDs are consistent with the combined median SED. Moreover, we verify that stacking the same exact 22 objects having NIRCam/MIRI coverage in all bands (from the GOODS-S field), the resulting SED is fully consistent with that obtained via stacking the full LRD sample. These checks strengthen that similarly-selected LRDs in different fields and with heterogeneous spectral coverage share a common median SED.
B.3 Effect of photometric redshifts
A relevant caveat in our analysis is the photometric selection of LRDs. Specifically, photometric redshift uncertainties can be notably high, potentially biasing the LRD selection. We iterate that for the spectroscopic subsample (85/302 objects) there is good agreement between photometric and spectroscopic redshifts (Sect. 2). However, we test this directly by stacking NIRCam and MIRI images solely for the spectroscopic subsample. The third panel in Fig. 5 displays the corresponding median stacked SED (blue filled circles). Numbers are given in Table 1 for each field: when combining all fields there are between 21 and 47 objects in MIRI (except for f560w, f1280w and f2550w with 5 objects due to a smaller areal coverage) and 85 objects in NIRCam. Despite the S/N being smaller than that of the full sample, the two median stacked SEDs agree remarkably well with each other.
B.4 Effect of LRD selection
In the literature, LRDs are inevitably selected with distinct criteria, depending on the availability of multi-band NIRCam and MIRI imaging, and rest-frame optical spectra. While our LRD sample meets the same compactness and V-shape criteria across all fields, we compare our stacked SED against other LRD SEDs based on different selections. The fourth panel of Fig. 5 shows for comparison: the median stacked SED from Akins et al. (2024; green symbols); the corresponding galaxy and AGN best-fit LRD SED from Casey et al. (2024; dot-dashed lines); the ALMA upper limit from Casey et al. (2025; downward triangle) and the observed CAPER-z9 NIRSpec spectrum (at =9.288; Taylor et al. 2025) shifted to =6.2 (orange symbols).
Akins et al. (2024) selected 434 LRDs in the COSMOS-Web area (0.5 deg2) that simultaneously meet a compactness (f444w) cut, a very red colour cut (f444w/f277w) and a bad brown drarf fit, without explicitly requiring a “V-shaped” SED due to limitation in NIRCam imaging. Their median SED agrees remarkably well with ours, though their statistics in MIRI is limited to 7.7m (over 1/3 of COSMOS-Web) and 18m (from the inner PRIMER-COSMOS footprint, 4% of the full sample), this latter leading to tentative detection (0.430.15 mJy). We show that their 18m flux is fully consistent with the rising MIR slope seen longward 10m. Their upper limits are here re-scaled to 3 (Herschel and SCUBA limits are too shallow to enter the y-scale). Despite the different LRD selection and survey field (25% of our LRDs are in Akins et al. 2024), this agreement further suggests that the compactness and the conservative colour selection by Akins et al. (2024) broadly compensate the need for a V-shaped SED, leading to self-consistent LRD samples.
In the same fourth panel of Fig. 5, we compare our stacks against the median LRD SED by Casey et al. (2024), that is inferred from a joint LRD sample (675 LRDs) from Akins et al. (2024) and Kokorev et al. (2024). This latter consists of a compactness (in f444w) criterion and various colour cuts that mimic a V-shape requirement (see their Sect.3.1 in Kokorev et al. 2024 for details). In their work, Casey et al. (2024) performed stacking in rest-frame wavelength bins, following the observed SED of each LRD. The median stacked SED was then redshifted to =6.2 (as for our sample) and fitted with either a pure AGN (red) or a pure galaxy template (blue). We emphasize that our median stacks are fully consistent with their median AGN SED within their uncertainties (here not reported). If any, the AGN SED from Casey et al. (2024) stands slightly higher than our MIRI points, that partly reflects the MIRI upper limits in their study (20 objects per bin at 0.8m). Our stacks and their SED are also mutually consistent with the ALMA upper limit derived in Casey et al. (2025) by stacking 60 LRDs (among the Akins et al. 2024 sample) at =1.3 mm.
In addition, we compare our stacks against the NIRSpec spectrum of CAPER-z9 at =9.288 (Taylor et al. 2025), which is complemented by MIRI data at 10m (detection) and 21m (upper limit). This continuous spectrum enables a direct comparison with our SED shape over a broad wavelength range. After moving the CAPER-z9 spectrum and photometry to =6.2 (orange symbols), we again find a reassuring consistency with our data.
Another meaningful comparison with MIRI studies of LRDs is against the sample of Williams et al. (2024), who stacked LRDs in all MIRI bands within the GOODS-SMILES footprint. They found a flat IR-SED up to 18m (and upper limits longward), thus excluding a dominant contribution from AGN-heated dust. Besides the small numbers (only 9 objects) and the slightly higher redshift compared to our sample (7.7 instead of 6.2), their LRDs were pre-selected from a pool of H-band dropouts (i.e. HST-dark galaxies), hence likely biased towards the most obscured sources. We repeat the stacking for the same 9 sources, finding the same median SED as in Williams et al. (2024). This check suggests that their sample might be sensitive to more obscured sources than conventional LRDs, especially considering their H-dropout selection, combined with the lack of compactness and V-shaped requirements. These differences might link to a different obscuration origin relative to typical LRDs. Hence we caution that these 9 LRDs are not representative of our photometrically-selected LRDs (from Kocevski et al. 2025), albeit they are certainly an interesting sub-sample to follow-up.
B.5 Effect of a broad redshift range
As shown in Fig. 4, our targets span a very broad redshift range, from 2 to 11, with over 90% of the sample being within the range (48). Because we perform stacking in the observed frame, the contribution to each stacked point comes from very different parts of the rest-frame spectrum. As mentioned in Sect. 4.1, we already account for this by convolving the filter response with the underlying redshift distribution. However, this approach assumes that each galaxy contributes evenly to the stack, regardless of its redshift, whereas the flux dimming can vary significantly across the full range. We thus test the impact of the redshift range to the stacked fluxes by splitting our LRDs into two similarly-populated -bins at 46 (126 LRDs) and 68 (121 LRDs). We show this comparison in the bottom panel of Fig. 5. The lower- bin shows slightly brighter fluxes and steeper MIRI slope compared to the higher- bin, as expected given the smaller distance and the longer rest-frame wavelength probed (e.g. 3.5 m instead of 2.5 m at MIRI-21 m), respectively. Overall, this check suggests that LRDs selected at different redshifts have broadly-consistent stacked SEDs, with no evidence for a galaxy-like drop at rest 1–2 m. We further note that emission lines are unlikely to systematically boost the median stacks, as this would be the case if a large fraction of sources were clustered in a narrow redshift range (such that a rest-frame emission line would systematically fall into a given filter). Given the broad similarity of the two SEDs in two different -bins, we believe this is unlikely the case.
Appendix C Input SED-fitting parameters
In the following Table, we outline the main input parameters used for SED-fitting with CIGALE (Section 4.1).
| Module | Parameter | Symbol | Values |
[…]
| Star formation history | Stellar age | ( yr) | 200, 300, 500, 700, 900∗ |
| Age of the burst | ( yr) | 10, 30, 50∗, 100 | |
| SFR ratio after/before the burst | 0.1, 0.5, 1∗, 2, 5, 10 ,30, 50 | ||
[…]
| (Bruzual & Charlot 2003) | Metallicity | 0.0004, 0.008, 0.02∗ | |
[…]
| Gas metallicity | Zgas | 0.0004∗, 0.008, 0.019 | |
| Electron density | 100,1000∗ | ||
[…]
| Dust attenuation | Fractional ISM attenuation | AV,ISM / (AV,BC+AV,ISM) | 0.44∗ |
| mod. (Charlot & Fall 2000) | Power-law slope (ISM) | – | -0.7∗ |
| Power-law slope (BC) | – | -1.3∗ | |
[…]
| Galactic dust emission | Minimum radiation field | Umin | 1.0, 10, 20, 40∗ |
| (Dale et al. 2014) | Slope in | 2.0∗ | |
| Illuminated fraction [Umin–Umax] | 0.02∗ | ||
[…]
| Slope of radial dust density profile | – | 1.0∗ | |
| Slope of polar dust density profile | – | 1.0∗ | |
| Torus opening angle (from equator) | 40∗ | ||
| AGN disk+torus+polar dust: | Ratio outer/inner torus radius | – | 20∗ |
| skirtor | Viewing angle (from vertical axis) | 50∗, 70 | |
| Stalevski et al. (2012; 2016) | AGN fraction | 0.0, 0.1, 0.3, 0.5, 0.7, 0.9, 0.99∗ | |
| Extinction law for polar dust | – | SMC | |
| in the polar direction | – | 0.00, 0.05∗, 0.10 | |
| Temperature of polar dust | – | 100 K∗ | |
| Emissivity of polar dust | – | 1.6∗ |
Appendix D Output SED-fitting parameters
| Parameter | Value and 1 error |
| Stellar mass (fit without AGN) | (5.01.0)109 M⊙ |
| Stellar mass (fit with AGN) | (2.50.3)109 M⊙ |
| Total dust luminosity (fit without AGN) | (5.41.2)1044 erg/s |
| Total dust luminosity (fit with AGN) | (1.40.2)1044 erg/s |
| SFR [100 Myr] (fit without AGN) | (8.23.6) M⊙/yr |
| SFR [100 Myr] (fit with AGN) | (0.60.1) M⊙/yr |
| SFR [10 Myr] (fit without AGN) | (15.95.2) M⊙/yr |
| SFR [10 Myr] (fit with AGN) | (0.50.2) M⊙/yr |
| AGN bolometric luminosity | (2.50.4)1044 erg/s |
| AGN 6 m luminosity | (3.80.6)1043 erg/s |

### v2

@@
 Little Red Dots (LRDs) are a puzzling population of extragalactic sources whose origin is highly debated.
-In this Letter, we stack NIRCam, MIRI and ALMA images of a large and homogeneously-selected sample of LRDs from multiple JWST Legacy fields.
-We find clear evidence for hot-dust emission, seen as a rising mid-infrared continuum up to rest-frame 3m, that is best explained by a standard dusty AGN structure.
-This scenario holds regardless of whether the Optical/Ultraviolet (UV) continuum is stellar or AGN-dominated.
-Either ways, we argue that AGN-heated dust must have a geometry that leaves at least part of the Optical/UV continuum and broad-line region emission unobscured, which explains the observed blue UV slopes and prominent Balmer features.
-Lack of detections in the deep stack of X-ray images suggests that Compton-thick (31024 cm-2) gas obscuration is common, and likely confined within the dust sublimation radius (sub0.1 pc).
-While this gas-dust displacement is in line with recent models (e.g.
-the “Black Hole Star”), how pre-enriched hot dust can form around a nearly-pristine black hole environment remains to be explained.
+In this work, we perform a comprehensive stacking analysis of NIRCam, MIRI and ALMA images of a large and homogeneously-selected sample of LRDs from multiple JWST Legacy fields.
+We report clear evidence for hot-dust emission in the median stacked spectral energy distribution (SED), featuring a rising near-infrared continuum up to rest-frame 3m, which is best explained by a standard dusty AGN structure.
+Although LRDs are likely a heterogeneous population, our findings suggest that most (50%) LRDs show AGN-heated dust emission, regardless of whether the Optical/Ultraviolet (UV) continuum is stellar or AGN-dominated.
+In either case, the best-fit dusty-AGN SED, combined with the lack of X-ray detection in the deep Chandra stacks, suggests that Compton-thick (31024 cm-2) gas obscuration is common, and likely confined within the dust sublimation radius (sub0.1 pc).
+Therefore, we argue that AGN-heated dust does not directly obscure either the Optical/UV continuum or the broad-line region emission, in order to explain the observed blue UV slopes and prominent Balmer features.
+While a gas-dust displacement is in line with several models, the formation scenario (in-situ or ex-situ) of this pre-enriched hot dust remains unclear.
 Key Words.:
 galaxies: nuclei – galaxies: active – galaxies: evolution1 Introduction
-The James Webb Space Telescope (JWST) has ushered extragalactic astronomy in a transformative era, offering unprecedented access to the high-redshift universe.
-One of the most puzzling discoveries is a population of sources dubbed as “Little Red Dots” (LRDs, e.g.
+The James Webb Space Telescope (JWST) has ushered extragalactic astronomy in a transformative era, offering unprecedented access to the high-redshift (4) universe, with the discovery of previously missed Active Galactic Nuclei (AGN; e.g.
+Matthee et al.
+2024).
+These AGNs are mostly identified via broad emission lines (FWHM1000–2000 km s-1), corresponding to bolometric luminosities Lbol,AGN1044-46 erg s-1 and harboring supermassive black holes (SMBHs, with mass ).
+Conversely to their lower- analogs, they do not show significant X-ray or radio emission (Ananna et al.
+2024, Mazzolari et al.
+2025).
+A non-negligible fraction of these AGN (10–30%; Hainline et al.
+2025) feature unique and puzzling properties, and are dubbed as “Little Red Dots” (LRDs, e.g.
 Matthee et al.
@@
 2025).
-LRDs are extremely compact (often unresolved at 100 pc scales) and show red optical colors combined with blue UV slopes, resulting in a characteristic “V-shaped” spectral energy distribution (SED), with a sharp turnover near the Balmer break at rest Å (e.g.
+These LRDs are extremely compact (often unresolved at 100 pc scales) and show red optical colors combined with blue Ultraviolet (UV) slopes, resulting in a characteristic “V-shaped” spectral energy distribution (SED), with a sharp turnover near the Balmer break at rest Å (e.g.
 Setton et al.
@@
 2025).
-Remarkably, LRDs appear to be two orders of magnitude more abundant than luminous quasars at similar redshifts (Kokorev et al.
+The vast majority of them are seen at 48, and about 60% of the spectroscopically-confirmed LRDs display broad emission lines (e.g.
+Greene et al.
+2024; Hviding et al.
+2025).
+Notably, LRDs appear to be up to two orders of magnitude more abundant than luminous quasars at z4 (Kokorev et al.
 2024, Akins et al.
@@
 However, an observational consensus of the physical nature of LRDs is still lacking.
-The red optical colors and broad emission lines point toward obscured Active Galactic Nuclei (AGN) activity (Inayoshi et al.
+On the one hand, if LRDs are compact star-forming galaxies (e.g., Labbé et al.
+2023), their stellar masses (108-11 M⊙) would imply extreme stellar densities (Pacucci & Narayan 2024).
+On the other hand, the red optical colors and broad emission lines hint at an AGN origin (Inayoshi et al.
 2020, Volonteri et al.
 2021, Ji et al.
+2025), from which the (e.g.
+Hα-based) estimates would imply overmassive BHs (/1:100; see e.g.
+Pacucci et al.
+2023; Juodžbalis et al.
+2024b; Maiolino et al.
+2024; Juodžbalis et al.
+2025) compared to local scaling relations (Reines & Volonteri 2015), even without accounting for dust attenuation (e.g.
+Greene et al.
 2025).
-In this case, the AGN continuum would be moderately reddened by dust (– mag; Casey et al.
+In the AGN scenario, the Optical/UV continuum would be moderately reddened by dust (– mag; Casey et al.
 2024; Akins et al.
@@
 2025).
-To alleviate this tension, some theoretical models postulate LRDs as supermassive black holes (SMBHs) accreting at super-Eddington luminosities and enshrouded in a turbulent, dust-free, and highly dense atmosphere: the “BH-star” scenario.
-This explains the strong Balmer break and the red continuum through pure gas opacity, and the broad Balmer lines via resonant electron scattering (e.g.
-Naidu et al.
-2025; Inayoshi & Maiolino 2025; Rusakov et al.
+Recent studies attribute the X-ray weakness to super-Eddington accretion (Madau & Haardt 2024; Lambrides et al.
+2024; Pacucci & Narayan 2024; King 2025; Madau 2025; but see e.g.
+Sacchi & Bogdán 2025).
+To also alleviate the tension between the prominent Optical/UV Balmer features and the apparent X-ray weakness, several models postulate the presence of a turbulent, dust-free, and highly dense gas envelope (e.g.
+Liu et al.
+2025; Cenci & Habouzit 2025; Inayoshi & Maiolino 2025; Naidu et al.
+2025; Rusakov et al.
 2025; de Graaff et al.
 2025).
-Similar gas conditions are postulated in other models, with the presence of a metal-free halo surrounding a pre-collapse “super-massive star” (Nandal & Loeb 2025), or a late-stage “Quasi star” (Begelman & Dexter 2025).
-All these models share the absence of prior metal enrichment and the presence of a gas-dense, dust-free region at sub-pc scales.
-So far, this dust paucity is supported by the flat LRD SED observed beyond rest 1.6 m (e.g.
-Williams et al.
-2024; Carranza-Escudero et al.
-2025), in contrast to the rising mid-infrared (MIR) continuum expected in obscured AGN, conveying the idea that MIR emission in LRDs is dominated by dust-poor stellar populations and not AGN dust (Williams et al.
-2024; Pérez-González et al.
-2024, Leung et al.
-2024).
-In addition, the non-detection from deep ALMA stacking poses a stringent dust mass limit of Mdust106 M⊙ at 6 (Casey et al.
+These conditions explaining the strong Balmer breaks and the red continua through pure gas opacity, since the gas temperature is too high for the dust to survive.
+In addition, the broad Balmer lines and other abrorption features can be explained via electron scattering (e.g.
+Rusakov et al.
+2025; Inayoshi et al.
+2025) or resonant Balmer scattering (e.g.
+Naidu et al.
+2025).
+These models mostly differ on the nature of the central object within the gas envelope, which could be either a direct-collapse BH (e.g.
+Cenci & Habouzit 2025), or a Super-Eddington SMBH (Liu et al.
+2025), or a late-stage “Quasi star” (Begelman & Dexter 2025), or a “BH-Star” (e.g.
+Naidu et al.
+2025), or other metal-free objects like a “super-massive star” (Nandal & Loeb 2025).
+Such a dust paucity is further supported by the non-detection from deep ALMA stacking, which poses a stringent upper limit to the cold dust budget, corresponding to Mdust106 M⊙ at 6 (Casey et al.
 2025; see also Chen et al.
@@
 2025).
-Nevertheless, we caution that previous NIRCam/MIRI SED studies of LRDs might have been biased towards bright objects (e.g.
+In the rest-frame near and mid-IR, a broad diversity of LRD SEDs has been reported in the recent literature.
+Stacking of MIRI-undetected LRDs reveals a flat SED beyond rest 1.6 m (e.g.
+Williams et al.
+2024; Carranza-Escudero et al.
+2025), in contrast to the rising mid-infrared (MIR) continuum expected in obscured AGN, conveying the idea that MIR emission in LRDs is typically dominated by dust-poor stellar populations and not AGN dust.
+However, based on individual MIRI detections, other studies found that some LRDs display a flat or rising MIR continuum (Pérez-González et al.
+2024; Barro et al.
+2024; Leung et al.
+2024; de Graaff et al.
+2025; Ronayne et al.
+2025), which is best fitted via composite stellar and AGN templates, suggesting a heterogeneous origin of the IR emission of LRDs.
+In this highly debated framework, we caution that previous MIRI-SED studies of LRDs might have been biased towards bright individual objects (e.g.
 Labbe et al.
@@
 2024), whose SED could intrinsically differ from that of “typical” LRDs.
-To explore this possibility, in this Letter we stack a large and homogeneously-selected sample of LRDs with broad NIRCam/MIRI coverage.
+To explore this possibility, in this paper we present a comprehensive stacking analysis of a large and homogeneously-selected LRD sample with broad NIRCam, MIRI (and ALMA) coverage.
 Throughout, we assume a Chabrier (2003) IMF and a flat CDM cosmology with = 70 km/s/Mpc, , and .
-2 Photometric selection of LRDs
-We select LRDs from the latest compilation of Kocevski et al.
-(2025), consisting of 341 photometrically-selected LRDs in the redshift range 2–11, and retrieved from several JWST/NIRCam surveys.
-These LRDs were homogeneously selected to display compact morphologies and “V-shaped” continuum from spectral slope fitting between blueward and redward of the rest-frame 3645 Å (i.e.
-close to the Balmer break).
-These criteria enable a self-consistent search for LRD candidates with red optical and blue UV colors over a wide redshift interval (Hainline et al.
-2025), although we acknowledge that it does not guarantee an homogeneous physical selection of LRDs.
-We replace photometric redshifts () with NIRSpec-based spectroscopic redshifts () taken from the public Dawn JWST Archive (DJA; Brammer 2023b; de Graaff et al.
-2024; Heintz et al.
-2024), whenever of high-quality (grade=3, see archive).
-This increases the spectroscopic fraction from 11% (39/341) to 29% (100/341).
-The mutual agreement between and estimates is generally good, with only five sources showing an offset /(1+)0.2.
-Finally, we restrict ourselves to the NIRCam fields with (at least partial) MIRI coverage in two or more bands (including 18m and/or 21m; see Table 1), which yield the strongest IR-SED constraints.
-Our final sample counts 302 LRDs taken from the following surveys: the Cosmic Evolution Early Release Science Survey (CEERS; Finkelstein et al.
-2023) in the EGS; the JWST Advanced Deep Extragalactic Survey (JADES; Eisenstein et al.
-2023) in the GOODS-S, complemented by MIRI imaging from the Systematic Mid-infrared Instrument Legacy Extragalactic Survey (SMILES; Rieke et al.
-2024; Alberts et al.
-2024); the Public Release IMaging for Extragalactic Research (PRIMER111https://primer-jwst.github.io/) survey in the COSMOS and UDS fields.
-Out of 302 LRDs, 163 (54%) are covered by MIRI 7.7m, 126 (42%) also by MIRI-18m, and 85 (28%) have a robust .
-Moreover, we exploit archival ALMA/B6 (1.1 mm) data from the combined A3COSMOS+A3GOODSS surveys (Liu et al.
-2019; Adscheid et al.
-2024; Magnelli et al.
-2024), which are available for a total of 70 LRDs in the COSMOS and GOODS-S fields.
-The number of sources split down by field, photometric coverage and spectroscopic redshifts is summarized in Table 1.
-3 Stacking results
-We perform median stacking over all available NIRCam bands (7) and all MIRI bands (2 in PRIMER, 7 in CEERS and 8 in SMILES), whenever available, at the optical position of all LRDs.
-We use a fixed pixel scale of 0.06′′, and we extract fluxes from a circular aperture with fixed radius =0.25′′.
-ALMA mean stacking, instead, is performed in the -plane, in order to homogeneously combine archival data at different frequency, angular resolution and sensitivity.
-Further details of the stacking methodology are given in Appendix A.
-We refer the reader to Appendix B, to verify the robustness of our stacking results against possible biases.
-In Fig.
-1 we show the stacked cutouts (3′′3′′) in each band, sorted by increasing wavelength.
-For comparison, the dashed white circle marks the corresponding PSF FWHM.
-As MIRI images are strongly oversampled, to aid visualization we smooth them with a Gaussian kernel, adopting a larger radius for larger PSF FWHM (see caption of Fig.
-1).
-The stacked ALMA/B6 image is obtained in the -plane and imaged at 0.1′′/px scale.
-Except for MIRI/f2550w and ALMA/B6, all other median stacks lead to a detection with signal-to-noise ratio (S/N) greater than 3.
-We stress that all detections are consistent with being unresolved, as their extent is smaller or comparable with the PSF FWHM of the corresponding band (white circles in Fig.
-1).
-All fluxes, errors, rms of the stacked images are listed in Table 2 along with the effective number of sources stacked in each band.
-4 Discussion and Conclusions
-4.1 SED-fitting with vs without AGN
-In order to interpret the stacked photometry of LRDs, we perform SED-fitting using the Code Investigating GALaxy Emission (CIGALE; Boquien et al.
-2019).
-We fix the input redshift to the median value of the LRD sample (6.2).
-In order to account for the broad redshift distribution of the sample (mostly at 4.58; see Fig.
-4) and the correspondingly wide rest-frame range sampled by each filter, we convolve the filter response curves at 6.2 with the band-specific redshift distribution of the underlying LRD sample.
-Then, we run CIGALE both with and without AGN templates to test the relative significance of the AGN component.
-In both SED runs, the library of stellar templates is taken from Bruzual & Charlot (2003), including a delayed star formation history with random bursts.
-Energy balance is assumed between the dust-absorbed UV radiation and the reprocessed IR emission, modeled via a dust attenuation curve from Charlot & Fall (2000), and dust templates from Dale et al.
-(2014).
-The AGN templates come from the skirtor module (Stalevski et al.
-2012; 2016), which includes emission from an accretion disk, a dusty torus and polar dust.
-A comprehensive list of the SED-fitting parameters is given in Appendix C and D.
-In Fig.
-2, we display all stacked fluxes (or 3 upper limits) overlaid to the total best-fit SED (black line) obtained from the run without AGN (top panel) or with AGN (middle panel).
-We caution the reader that our SED-fitting analysis is not primarily aimed at interpreting the rest optical or far-IR properties of LRDs.
-For instance, CIGALE assumes a single-zone obscuration between optical, UV and IR light, which might be oversimplistic in case of no co-spatiality (see Sect.
-4.2).
-Nonetheless, regardless of the origin of Optical/UV emission in LRDs, we argue that AGN-heated dust is needed to reproduce the monothonic increase up to rest-frame 2–3m, which pure dust-attenuated stellar light fails to match.
-This is further confirmed by our ALMA upper limit, which is very similar to that derived by Casey et al.
-(2025) (see also Xiao et al.
-2025; Setton et al.
-2025), which translates into Mdust106 M⊙ for T100 K.
-Indeed, the best-fit SED without AGN strongly overshoots the ALMA upper limit when attempting to reproduce the power-law-like SED at rest 1–3m.
-While this could be resolved by lifting the energy balance implemented in CIGALE, even ignoring the upper limits (at f2550w and ALMA/B6), a simple F-test comparing the two reduced values favours the AGN solution at 95% significance.
-4.2 Are LRDs typically dust-obscured AGN?
-By considering the best-fit solution with AGN, the rest-frame 6 m luminosity of the AGN torus can be used as a proxy for the intrinsic AGN X-ray (rest [2–10] keV) luminosity (X,intr (e.g.
-Stern 2015).
-From 6μm=(3.80.6)1043 erg/s, we derive X,intr=21043 erg/s, which nicely matches that expected from the best-fit AGN bolometric luminosity, if assuming standard X-ray bolometric corrections (e.g.
-Lusso et al.
-2012).
-By stacking the deepest X-ray images of the 44 LRDs falling inside the Chandra Deep Field South (Luo et al.
-2017), we obtain a 3 [2–10] keV luminosity limit X1042 erg/s at =6.2 (Comastri et al.
-subm).
-In order to match it with the X,intr expected from the AGN SED, the hydrogen column density has to be NH31024 cm-2, implying Compton-thick gas obscuration (Ananna et al.
-2024; Yue et al.
-2024; Maiolino et al.
-2025).
-Given the SED-based estimate of the AGN bolometric luminosity, erg s-1 and the best-fit stellar mass from the AGN solution (M⋆2.5109 M⊙; i.e.
-2 smaller than from a pure-galaxy fit), we emphasize that, if LRDs are highly-accreting AGN (Eddington ratio 10%), the BH-to-galaxy mass ratio MBH/M⋆1:100, implying strongly overmassive BHs compared to local scaling relations (see also Pacucci et al.
-2023; Juodžbalis et al.
-2024; Maiolino et al.
-2024).
-Moreover, our estimate corresponds to a dust sublimation radius of 0.1 pc at K (Barvainis 1987), meaning that any dusty AGN structure should begin at scales pc from the central SMBH.
-In the bottom panel of Fig.
-2, we try to decompose the AGN-dust template with three multi-temperature greybody curves (with free to vary).
-Our rising MIRI SED up to rest-frame 2–3m indicates hot dust at K, though it could be even hotter if optically-thick.
-Instead, the bulk dust mass (Mdust3104 M⊙) comes from colder polar AGN dust at 50 K (though not constrained by our data).
-While a dusty structure could induce X-ray obscuration, it would also strongly suppress the Optical/UV continuum, in tension with the blue UV slope and strong Balmer features that imply relatively modest dust attenuation (; Akins et al.
-2024; Casey et al.
-2025).
-This apparent contradiction can be resolved if X-ray absorption in LRDs occurs within , in a gas-dense and dust-free region.
-This gas-dust displacement is formally consistent with the recently-proposed models of Quasi-star (Begelman & Dexter 2025), BH–star (Naidu et al.
-2025), and super-massive star (Nandal & Loeb 2025) scenarios.
-However, the presence of pre-enriched dust around a first-generation (or metal-poor) star is hard to justify.
-Moreover, the normalized BH–star SED (green line in Fig.
-2, bottom panel) appears to flatten out at rest-frame 1 m, contrary to the rising MIRI fluxes.
-Our findings might fit better with an alternative hypothesis suggested by previous studies (Setton et al.
-2024; Maiolino et al.
-2025; D’Eugenio et al.
-2025; Ji et al.
-2025; Inayoshi & Maiolino 2025), in which X-ray obscuration arises from Compton-thick gas clouds within the BLR (at sub, i.e.
-dust-free), which covers the compact X-ray corona without suppressing the broad Balmer lines.
-Yet, the origin of the Optical/UV continuum remains open.
-Although we caution that LRDs might not consist of a single population, we speculate two simple scenarios in Fig.
-3, depending on what powers the Optical/UV continuum in LRDs.
-In either cases, the bulk of Optical/UV continuum is not dust-attenuated, despite the presence of hot AGN dust, and the outer BLR is directly visible along the line of sight.
-(i) If the Optical/UV comes from stellar light (Fig.
-3, left), the LRD resembles a “type-2” AGN configuration, in which the torus covers the AGN accretion disk, which would otherwise outshine the host.
-In case of high dust clumpiness, mildly-attenuated light from the outer accretion disk could add to the optical emission.
-(ii) Instead, if the Optical/UV comes from AGN light (Fig.
-3, right), the LRD resembles a “type-1” unobscured AGN configuration, though the origin of the sharp Balmer break remains unclear.
-Fainter Optical/UV light from kpc-scale star formation cannot be ruled out (e.g.
-Rinaldi et al.
-2025).
-In summary, we find clear evidence for a rising MIRI continuum up to 3m, which we interpret via SED fitting as hot (8205 K) AGN dust, regardless of whether the Optical/UV continuum is stellar- or AGN-dominated, or a combination thereof.
-Despite the presence of hot dust, these LRDs selected via V-shaped and compactness criteria must be not strongly dust-obscured.
-We acknowledge that such homogeneous selection does not necessarily translate into a single LRD population with common physical properties.
-Our findings mainly provide en empirical benchmark for the presence of AGN-heated dust in the majority of LRDs, or (at least) in a significant sub-population of them.
-For instance, if LRDs are standard AGN, we speculate that the above selection criteria might be biased against the most dust-obscured (“type-2”) LRD analogs, which could be recovered by lifting the blue-UV slope requirement.
-Acknowledgements.
-ID acknowledges funding by the European Union – NextGenerationEU, RRF M4C2 1.1, Project 2022JZJBHM: ”AGN-sCAN: zooming-in on the AGN-galaxy connection since the cosmic noon” - CUP C53D23001120006.
-(Some of) the data producs presented here were retrieved from the DJA, an initiative of the Cosmic Dawn Center (DAWN), which is funded by the Danish National Research Foundation under grant DNRF140.References
-- Adscheid et al. (2024) Adscheid, S., Magnelli, B., Liu, D., et al. 2024, A&A, 685, A1
-- Akins et al. (2024) Akins, H. B., Casey, C. M., Lambrides, E., et al. 2024, arXiv e-prints, arXiv:2406.10341
-- Alberts et al. (2024) Alberts, S., Lyu, J., Shivaei, I., et al. 2024, ApJ, 976, 224
-- Ananna et al. (2024) Ananna, T. T., Bogdán, Á., Kovács, O. E., Natarajan, P., & Hickox, R. C. 2024, ApJ, 969, L18
-- Barro et al. (2024) Barro, G., Perez-Gonzalez, P. G., Kocevski, D. D., et al. 2024, arXiv e-prints, arXiv:2412.01887
-- Barvainis (1987) Barvainis, R. 1987, ApJ, 320, 537
-- Begelman & Dexter (2025) Begelman, M. C. & Dexter, J. 2025, arXiv e-prints, arXiv:2507.09085
-- Boquien et al. (2019) Boquien, M., Burgarella, D., Roehlly, Y., et al. 2019, A&A, 622, A103
-- Brammer (2023a) Brammer, G. 2023a
-- Brammer (2023b) Brammer, G. 2023b
-- Bruzual & Charlot (2003) Bruzual, G. & Charlot, S. 2003, MNRAS, 344, 1000
-- Carranza-Escudero et al. (2025) Carranza-Escudero, M., Conselice, C. J., Adams, N., et al. 2025, arXiv e-prints, arXiv:2506.04004
-- Casey et al. (2025) Casey, C. M., Akins, H. B., Finkelstein, S. L., et al. 2025, arXiv e-prints, arXiv:2505.18873
-- Casey et al. (2024) Casey, C. M., Akins, H. B., Kokorev, V., et al. 2024, ApJ, 975, L4
-- Chabrier (2003) Chabrier, G. 2003, PASP, 115, 763
-- Charlot & Fall (2000) Charlot, S. & Fall, S. M. 2000, ApJ, 539, 718
-- Chen et al. (2025) Chen, K., Li, Z., Inayoshi, K., & Ho, L. C. 2025, arXiv e-prints, arXiv:2505.22600
-- Dale et al. (2014) Dale, D. A., Helou, G., Magdis, G. E., et al. 2014, ApJ, 784, 83
-- de Graaff et al. (2024) de Graaff, A., Rix, H.-W., Carniani, S., et al. 2024, A&A, 684, A87
-- de Graaff et al. (2025) de Graaff, A., Rix, H.-W., Naidu, R. P., et al. 2025, arXiv e-prints, arXiv:2503.16600
-- D’Eugenio et al. (2025) D’Eugenio, F., Maiolino, R., Perna, M., et al. 2025, arXiv e-prints, arXiv:2503.11752
-- Eisenstein et al. (2023) Eisenstein, D. J., Johnson, B. D., Robertson, B., et al. 2023, arXiv e-prints, arXiv:2310.12340
-- Finkelstein et al. (2023) Finkelstein, S. L., Bagley, M. B., Ferguson, H. C., et al. 2023, ApJ, 946, L13
-- Franco et al. (2025) Franco, M., Casey, C. M., Koekemoer, A. M., et al. 2025, arXiv e-prints, arXiv:2506.03256
-- Gloudemans et al. (2025) Gloudemans, A. J., Duncan, K. J., Eilers, A.-C., et al. 2025, ApJ, 986, 130
-- Greene et al. (2024) Greene, J. E., Labbe, I., Goulding, A. D., et al. 2024, ApJ, 964, 39
-- Hainline et al. (2025) Hainline, K. N., Maiolino, R., Juodžbalis, I., et al. 2025, ApJ, 979, 138
-- Harish et al. (2025) Harish, S., Kartaltepe, J. S., Liu, D., et al. 2025, arXiv e-prints, arXiv:2506.03306
-- Heintz et al. (2024) Heintz, K. E., Watson, D., Brammer, G., et al. 2024, Science, 384, 890
-- Inayoshi & Maiolino (2025) Inayoshi, K. & Maiolino, R. 2025, ApJ, 980, L27
-- Inayoshi et al. (2020) Inayoshi, K., Visbal, E., & Haiman, Z. 2020, ARA&A, 58, 27
-- Ji et al. (2025) Ji, X., Maiolino, R., Übler, H., et al. 2025, arXiv e-prints, arXiv:2501.13082
-- Juodžbalis et al. (2024) Juodžbalis, I., Maiolino, R., Baker, W. M., et al. 2024, Nature, 636, 594
-- Kocevski et al. (2025) Kocevski, D. D., Finkelstein, S. L., Barro, G., et al. 2025, ApJ, 986, 126
-- Kokorev et al. (2024) Kokorev, V., Caputi, K. I., Greene, J. E., et al. 2024, ApJ, 968, 38
-- Labbe et al. (2025) Labbe, I., Greene, J. E., Bezanson, R., et al. 2025, ApJ, 978, 92
-- Labbe et al. (2024) Labbe, I., Greene, J. E., Matthee, J., et al. 2024, arXiv e-prints, arXiv:2412.04557
-- Leung et al. (2024) Leung, G. C. K., Finkelstein, S. L., Pérez-González, P. G., et al. 2024, arXiv e-prints, arXiv:2411.12005
-- Liu et al. (2019) Liu, D., Schinnerer, E., Groves, B., et al. 2019, ApJ, 887, 235
-- Luo et al. (2017) Luo, B., Brandt, W. N., Xue, Y. Q., et al. 2017, ApJS, 228, 2
-- Lusso et al. (2012) Lusso, E., Comastri, A., Simmons, B. D., et al. 2012, MNRAS, 425, 623
-- Magnelli et al. (2024) Magnelli, B., Adscheid, S., Wang, T.-M., et al. 2024, A&A, 688, A55
-- Maiolino et al. (2025) Maiolino, R., Risaliti, G., Signorini, M., et al. 2025, MNRAS, 538, 1921
-- Maiolino et al. (2024) Maiolino, R., Scholtz, J., Curtis-Lake, E., et al. 2024, A&A, 691, A145
-- Matthee et al. (2024) Matthee, J., Naidu, R. P., Brammer, G., et al. 2024, ApJ, 963, 129
-- Naidu et al. (2025) Naidu, R. P., Matthee, J., Katz, H., et al. 2025, arXiv e-prints, arXiv:2503.16596
-- Nandal & Loeb (2025) Nandal, D. & Loeb, A. 2025, arXiv e-prints, arXiv:2507.12618
-- Pacucci et al. (2023) Pacucci, F., Nguyen, B., Carniani, S., Maiolino, R., & Fan, X. 2023, ApJ, 957, L3
-- Pérez-González et al. (2024) Pérez-González, P. G., Barro, G., Rieke, G. H., et al. 2024, ApJ, 968, 4
-- Rieke et al. (2024) Rieke, G. H., Alberts, S., Shivaei, I., et al. 2024, ApJ, 975, 83
-- Rinaldi et al. (2025) Rinaldi, P., Rieke, G. H., Wu, Z., et al. 2025, arXiv e-prints, arXiv:2507.17738
-- Rusakov et al. (2025) Rusakov, V., Watson, D., Nikopoulos, G. P., et al. 2025, arXiv e-prints, arXiv:2503.16595
-- Setton et al. (2024) Setton, D. J., Greene, J. E., de Graaff, A., et al. 2024, arXiv e-prints, arXiv:2411.03424
-- Setton et al. (2025) Setton, D. J., Greene, J. E., Spilker, J. S., et al. 2025, arXiv e-prints, arXiv:2503.02059
-- Stalevski et al. (2012) Stalevski, M., Fritz, J., Baes, M., Nakos, T., & Popović, L. Č. 2012, MNRAS, 420, 2756
-- Stalevski et al. (2016) Stalevski, M., Ricci, C., Ueda, Y., et al. 2016, MNRAS, 458, 2288
-- Stern (2015) Stern, D. 2015, ApJ, 807, 129
-- Taylor et al. (2025) Taylor, A. J., Kokorev, V., Kocevski, D. D., et al. 2025, ApJ, 989, L7
-- Valentino et al. (2023) Valentino, F., Brammer, G., Gould, K. M. L., et al. 2023, ApJ, 947, 20
-- Volonteri et al. (2021) Volonteri, M., Habouzit, M., & Colpi, M. 2021, Nature Reviews Physics, 3, 732
-- Wang et al. (2025) Wang, B., de Graaff, A., Davies, R. L., et al. 2025, ApJ, 984, 121
-- Wang et al. (2024) Wang, T.-M., Magnelli, B., Schinnerer, E., et al. 2024, A&A, 681, A110
-- Wang et al. (2022) Wang, T.-M., Magnelli, B., Schinnerer, E., et al. 2022, A&A, 660, A142
-- Williams et al. (2024) Williams, C. C., Alberts, S., Ji, Z., et al. 2024, ApJ, 968, 34
-- Xiao et al. (2025) Xiao, M., Oesch, P. A., Bing, L., et al. 2025, arXiv e-prints, arXiv:2503.01945
-- Yue et al. (2024) Yue, M., Eilers, A.-C., Ananna, T. T., et al. 2024, ApJ, 974, L26
-Appendix A Stacked samples and fluxes
-Here we provide a more detailed outline of the input images and stacking technique used in this work.
-Image and noise maps are retrieved from the corresponding mosaics of each field (see below).
-In this step, we discard narrow or medium-band filters (whenever available) to mitigate potential flux boosting in the case of strong emission lines.
-We retrieve PRIMER-COSMOS images as part of the COSMOS-Web team, where the full PRIMER-COSMOS mosaic has been reduced following the same procedure as with the COSMOS-Web data (Franco et al.
-2025; Harish et al.
-2025).
-All NIRCam and MIRI mosaics have a pixel size of 0.06′′.
-For the GOODS-S, we retrieve NIRCam and MIRI images from the JADES and SMILES surveys publicly available on the MAST222SMILES: https://archive.stsci.edu/hlsp/smiles; JADES: https://archive.stsci.edu/hlsp/jades archive.
-While SMILES maps are already at 0.06′′/px, JADES maps (with a native pixel scale of 0.03′′) are re-scaled to 0.06′′/px for consistency with the other fields.
-For CEERS and PRIMER-UDS, which are unavailable on MAST, we leverage the DJA (Brammer 2023a; Valentino et al.
-2023), where the full reduced mosaics are publicly available in each NIRCam and MIRI band333DJA full reduced mosaics in each band are taken from https://s3.amazonaws.com/grizli-v2/JwstMosaics/v7/index.html, all with a pixel scale of 0.04′′ and in units of [10 nJy].
-We do a resampling of all mosaics to 0.06′′/px for consistency with SMILES and PRIMER-COSMOS maps, while preserving astrometry and total flux over a given aperture.
-For PRIMER-UDS, we also merge the East and West mosaics (listed separately in the DJA) into a full mosaic.
-We only consider sources whose position in the map is at least 20 pixels inside the edge of the mosaic.
-For each object and in each band, we create a pixel image (with =51, 0.06′′/px) from the corresponding map, each centred on the NIRCam position of the LRD.
-We then stack all stack sources and retrieve the median flux at each pixel.
-The uncertainty on the stacked flux density is quantified through a bootstrapping technique: being stack the number of input targets in a given band, in each random realisation we re-shuffle the input sample, preserving the same stack by allowing source duplication.
-We repeat this 100 times, and take the median of the resulting flux distribution as our formal stacked flux.
-The 1 dispersion around this value is interpreted as the flux error.
-The rms of the stacked image is computed by stacking the inverse-variance weighted noise map of each target.
-Our median stacks yield a detection in each band up to 21 m with signal-to-noise444We consider as “noise” the nominal rms of the stacked image; though using the flux error derived via bootstrapping does not change our conclusions.
-S/N3 (see Table 2).
-For ALMA, we perform a stacking analysis in the uv-domain using all publicly available Band-6 observations covering our LRDs in the COSMOS and GOODS-S fields.
-This analysis is carried out with CASA, following the method described in Wang et al.
-(2022) (see also Wang et al.
-2024; Magnelli et al.
-2024), to which we refer for further details of the procedure.
-Fig.
-4 displays the redshift distribution of our final LRD sample (302 sources, grey dashed), split by field (coloured open histograms).
-Filled histograms highlight the spectroscopic subsets.
-Corresponding numbers are listed in Table 1.
 | Survey/Field | Instrument | Filters | Area | stack range |
@@
 | SMILES/GOODSS | MIRI | [f560w, f770w, f1000w, f1280w, | ||
-| f1500w, f1800w, f2100w, f2550w] | 34.0 | 22–23 (2) | ||
+| f1500w, f1800w, f2100w, f2550w] | 34.0 | 22 (2) | ||
 | A3-GOODSS | ALMA | Band-6 (1080m) | 86.8 | 34 (4) |
-| PRIMER/COSMOS | NIRCam | [f090w, f115w, f150w, f200w, f277w, f356w, f444w] | 138.9 | 79–80 (7) |
-| MIRI | [f770w, f1800w] | 73.0 | 41–42 (3) | |
+| PRIMER/COSMOS | NIRCam | [f090w, f115w, f150w, f200w, f277w, f356w, f444w] | 138.9 | 79 (15) |
+| MIRI | [f770w, f1800w] | 73.0 | 41 (7) | |
 | A3-COSMOS | ALMA | Band-6 (1080m) | 127.9 | 36 (3) |
 | PRIMER/UDS | NIRCam | [f090w, f115w, f150w, f200w, f277w, f356w, f444w] | 243.0 | 116 (42) |
-| MIRI | [f770w, f1800w] | 130.0 | 61–62 (23) |
+| MIRI | [f770w, f1800w] | 130.0 | 61 (23) | |
+| All fields | NIRCam | – | 531.7 | 302 (96) |
+| MIRI | – | 292.0 | 163 (51) |
+2 Photometric selection of LRDs
+We select LRDs from the compilation of Kocevski et al.
+(2025), consisting of 341 photometrically-selected LRDs in the redshift range 2–11, and retrieved from several JWST/NIRCam surveys.
+These LRDs were homogeneously selected to display compact morphologies and “V-shaped” continua from spectral slope fitting between blueward and redward of the Balmer break (rest-frame 3645 Å).
+These criteria enable a self-consistent search for LRD candidates with red optical and blue UV colors over a wide redshift interval (Hainline et al.
+2025), although we acknowledge that it does not guarantee an homogeneous physical selection of LRDs.
+We replace photometric redshifts () with NIRSpec-based spectroscopic redshifts () taken from NIRSpec Merged Table (version 4222https://zenodo.org/records/15472354; Valentino et al.
+2025) of the public Dawn JWST Archive (DJA; Brammer 2023b; de Graaff et al.
+2024; Heintz et al.
+2025), whenever of high-quality (grade=3, see archive).
+We augment this list by adding measurements from the JADES public database333https://jades.herts.ac.uk/search/ (for the GOODS-S).
+This enhances the spectroscopic fraction from 11% (39/341) to 31% (107/341).
+The mutual agreement between and estimates is very good, with a median absolute deviation (MAD) 1.6%, and only five catastrophic outliers with /(1+)0.2.
+Hence, we do expect a negligible impact of photo- uncertainties on the median stacks.
+Fig.
+1 displays the redshift distribution of our final LRD sample (302 sources, grey dashed), split by field (coloured open histograms).
+Filled histograms highlight the spectroscopic subsets.
+Corresponding numbers are listed in Table 1.
+Finally, we restrict ourselves to the NIRCam fields with (at least partial) MIRI coverage in two or more bands (including 18m and/or 21m; see Table 1), which yield the strongest IR-SED constraints.
+Our final sample counts 302 LRDs taken from the following surveys: the Cosmic Evolution Early Release Science Survey (CEERS; Finkelstein et al.
+2023) in the EGS; the JWST Advanced Deep Extragalactic Survey (JADES; Eisenstein et al.
+2023) in the GOODS-S, complemented by MIRI imaging from the Systematic Mid-infrared Instrument Legacy Extragalactic Survey (SMILES; Rieke et al.
+2024; Alberts et al.
+2024); the Public Release IMaging for Extragalactic Research (PRIMER444https://primer-jwst.github.io/) survey in the COSMOS and UDS fields.
+Out of 302 LRDs, 163 (54%) are covered by MIRI 7.7m, 126 (42%) also by MIRI-18m, and 96 (32%) have a robust .
+Moreover, we exploit archival ALMA/B6 (1.1 mm) data from the combined A3COSMOS+A3GOODSS surveys (Liu et al.
+2019; Adscheid et al.
+2024; Magnelli et al.
+2024), which are available for a total of 70 LRDs in the COSMOS and GOODS-S fields.
+The number of sources split down by field, photometric coverage and spectroscopic redshifts is summarized in Table 1.
+3 Stacking method and results
+Here we outline the input images used for stacking and the stacking method adopted in this work.
+We perform median stacking in all available NIRCam bands (7) and MIRI bands (2 in PRIMER, 7 in CEERS and 8 in SMILES), at the optical position of all LRDs.
+We use a fixed pixel scale of 0.06′′, and we extract fluxes from a circular aperture with fixed radius =0.25′′.
+Image and noise maps are retrieved from the corresponding mosaics of each field (see below).
+In this step, we discard narrow or medium-band filters (whenever available) to mitigate potential flux boosting in the case of strong emission lines.
+We note that stacking is performed in the observed frame, since stacking in the rest-frame would require us to K-correct fluxes by assuming an intrinsic SED, which is instead what we would like to assess, particularly in the rest-frame near-IR (NIR, 1–3m), where an AGN-like or a galaxy-like SED alters the interpretation.
+However, a rest-frame stacking analysis was carried out in Casey et al.
+(2024), delivering consistent results with ours.
+We discuss this comparison in Sect.
+4.2.
+3.1 Input image maps
+We retrieve PRIMER-COSMOS images as part of the COSMOS-Web team, where the full PRIMER-COSMOS mosaic has been reduced following the same procedure as with the COSMOS-Web data (Franco et al.
+2025; Harish et al.
+2025).
+All NIRCam and MIRI mosaics have a pixel size of 0.06′′.
+For the GOODS-S, we retrieve NIRCam and MIRI images from the JADES and SMILES surveys publicly available on the MAST555SMILES: https://archive.stsci.edu/hlsp/smiles; JADES: https://archive.stsci.edu/hlsp/jades archive.
+While SMILES maps are already at 0.06′′/px, JADES maps are re-scaled from their native resolution (0.03′′/px) to 0.06′′/px for consistency with the other fields.
+For CEERS and PRIMER-UDS, which are unavailable on MAST, we leverage the DJA (Brammer 2023a; Valentino et al.
+2023), where the full reduced mosaics are publicly available in each NIRCam and MIRI band666DJA full reduced mosaics in each band are taken from https://s3.amazonaws.com/grizli-v2/JwstMosaics/v7/index.html, all with a pixel scale of 0.04′′ and in units of [10 nJy].
+We do a resampling of all mosaics to 0.06′′/px for consistency with SMILES and PRIMER-COSMOS maps, while preserving astrometry and total flux over a given aperture.
+For PRIMER-UDS, we also merge the East and West mosaics (listed separately in the DJA) into a full mosaic.
+We only consider sources whose position in the map is at least 20 pixels inside the edge of the mosaic.
+For each object and in each band, we create a pixel image (with =51, 0.06′′/px) from the corresponding map, each centred on the NIRCam position of the LRD.
+We then stack all stack sources and retrieve the median flux at each pixel.
+For ALMA, instead, we perform mean stacking in the -plane, using all publicly available Band-6 observations covering our LRDs in the COSMOS and GOODS-S fields.
+Mean stacking is needed in order to homogeneously combine archival data at different frequency, angular resolution and sensitivity.
+This analysis is carried out with CASA, following the method described in Wang et al.
+(2022) (see also Wang et al.
+2024; Magnelli et al.
+2024), to which we refer for further details of the procedure.
+3.2 Stacked fluxes
+In Fig.
+2 we show the stacked cutouts (3′′3′′) in each band, sorted by increasing wavelength.
+For comparison, the dashed white circle marks the corresponding PSF FWHM.
+As MIRI images are strongly oversampled, to aid visualization we smooth them with a Gaussian kernel, adopting a larger radius for larger PSF FWHM (see caption of Fig.
+2).
+The stacked ALMA/B6 image is obtained in the -plane and imaged at 0.1′′/px scale.
+Except for MIRI/f2550w and ALMA/B6, all other median stacks lead to a detection with signal-to-noise ratio (S/N) greater than 3.
+We stress that all detections are consistent with being unresolved, as their extent is smaller or comparable with the PSF FWHM of the corresponding band (white circles in Fig.
+2).
+The uncertainty on the stacked flux density is quantified through a bootstrapping technique: being stack the number of input targets in a given band, in each random realisation we re-shuffle the input sample, preserving the same stack by allowing source duplication.
+We repeat this 100 times, and take the median of the resulting flux distribution as our formal stacked flux.
+The 1 dispersion around this value is interpreted as the flux error.
+The rms of the stacked image is computed by stacking the inverse-variance weighted noise map of each target.
+Our median stacks yield a detection in each band up to 21 m with signal-to-noise777We consider as “noise” the nominal rms of the stacked image; though using the flux error derived via bootstrapping does not change our conclusions.
+S/N3.
+All fluxes, errors, rms of the stacked images are listed in Table 2 along with the effective number of sources stacked in each band.
+We refer the reader to Appendix A, where we test the robustness of our stacking results against possible biases.
+In particular, it is worth noting that (rms-weighted) mean stacking delivers systematically higher fluxes than median stacking, by a factor of 2–3 in all bands.
+In Appendix A.2, we demonstrate that, at least in MIRI bands, these fluctuations are entirely due to individual MIRI detections (dropping from 76% at f770w to 0% at f2100w), without which the mean and median stacked fluxes are matched one another.
+Therefore, we reiterate that mean stacking is highly sensitive to a few brighter detections, whereas median stacking is more representative of the bulk LRD population.
 | wavelength | stack | rms | Flux | Error Flux |
@@
 | 1080.00 | 70 | 5.89e-03 | ¡1.77e-02 | — |
-Appendix B Stacking tests
-We test our stacking results against multiple potential biases related to sample selection, stacking method, field-to-field variations and photometric redshift uncertainties.
-The main outcome of these tests is shown in Fig.
-5.
-In general, these tests further corroborate the robustness of our stacking results and the need for an obscured AGN component to reproduce the MIR excess.
-Each of the following tests is discussed in a dedicated subsection, numbered as the panels of Fig.
-5 from top to bottom.
-B.1 Mean vs median stacking
-In the first panel of Fig.
-5 we compare our median stacked fluxes (grey open circles) against mean stacked fluxes (red starred symbols).
-These latter values are computed from the inverse-variance weighted flux density map, accounting for the different sensitivity (i.e.
-exposure time) of each target.
-Mean stacking leads to higher fluxes with respect to median stacking.
-This is likely caused by the fact that the resulting (linear) mean flux can be boosted upwards by relatively bright individual detections (we note that upper limits at f2550w and ALMA/B6 are unchanged as they are calculated from the rms map).
-This is the case in NIRCam bands, where all LRDs are individually detected by definition over multiple bands, both blueward and redward the Balmer break (in order to assess the V-shape).
-A similar effect is seen at MIRI 7.7m and 18m, which include about 100 more targets than at other MIRI bands (see Table 1), thanks to a complete coverage across all the fields studied here.
-Even a handful of bright detections at 7.7m and 18m (some have S/N100 in the PRIMER fields) are able to boost the mean MIRI flux upwards.
-Moreover, we opt for median stacking also because it strongly mitigates contamination from bright neighbors and catastrophic outliers, and thus reduces the confusion noise for the faint sources.
-This comparison, however, demonstrates that the mean stacked LRD SED reveals a similar trend as the median stacked SED, showing a rising MIR slope that is indicative of dusty AGN activity.
-B.2 Field-to-field variations
-An important aspect that our analysis relies on is the combination of multiple NIRCam and MIRI survey fields.
-While this is necessary to improve the statistics and the stacked S/N, it might potentially introduce systematics related to different JWST coverage and depth, which could affect the properties of photometrically-selected LRDs.
-We thus verify that our stacking results are robust against field-to-field variations, by repeating the same analysis separately for each field.
-The second panel in Fig.
-5 displays the median stacked SED obtained in each field (represented by different filled symbols) overlaid to our full stacking results (grey open circles).
-For visual purposes, here we relax the detection threshold to S/N2 (hence upper limit fluxes are set to twice the noise level if S/N2).
-Despite their heterogeneous spectral coverage across MIRI bands, all single-field median SEDs are consistent with the combined median SED.
-Moreover, we verify that stacking the same exact 22 objects having NIRCam/MIRI coverage in all bands (from the GOODS-S field), the resulting SED is fully consistent with that obtained via stacking the full LRD sample.
-These checks strengthen that similarly-selected LRDs in different fields and with heterogeneous spectral coverage share a common median SED.
-B.3 Effect of photometric redshifts
-A relevant caveat in our analysis is the photometric selection of LRDs.
-Specifically, photometric redshift uncertainties can be notably high, potentially biasing the LRD selection.
-We iterate that for the spectroscopic subsample (85/302 objects) there is good agreement between photometric and spectroscopic redshifts (Sect.
-2).
-However, we test this directly by stacking NIRCam and MIRI images solely for the spectroscopic subsample.
-The third panel in Fig.
-5 displays the corresponding median stacked SED (blue filled circles).
-Numbers are given in Table 1 for each field: when combining all fields there are between 21 and 47 objects in MIRI (except for f560w, f1280w and f2550w with 5 objects due to a smaller areal coverage) and 85 objects in NIRCam.
-Despite the S/N being smaller than that of the full sample, the two median stacked SEDs agree remarkably well with each other.
-B.4 Effect of LRD selection
+4 Discussion and Conclusions
+4.1 SED-fitting with vs without AGN
+In order to interpret the stacked photometry of LRDs, we perform SED-fitting using the Code Investigating GALaxy Emission (CIGALE; Boquien et al.
+2019).
+We fix the input redshift to the median value of the LRD sample (6.2).
+In order to account for the broad redshift distribution of the sample (mostly at 4.58; see Fig.
+1) and the correspondingly wide rest-frame range sampled by each filter, we convolve the filter response curves at 6.2 with the band-specific redshift distribution of the underlying LRD sample.
+Then, we run CIGALE both with and without AGN templates to test the relative significance of the AGN component.
+In both SED runs, the library of stellar templates is taken from Bruzual & Charlot (2003), including a delayed star formation history with random bursts.
+Energy balance is assumed between the dust-absorbed UV radiation and the reprocessed IR emission, modeled via a dust attenuation curve from Charlot & Fall (2000), and dust templates from Dale et al.
+(2014).
+The AGN templates come from the skirtor module (Stalevski et al.
+2012; 2016), which includes emission from an accretion disk, a dusty torus and polar dust.
+A comprehensive list of the SED-fitting parameters is given in Appendix B and C.
+In Fig.
+3, we display all stacked fluxes (or 3 upper limits) overlaid to the total best-fit SED (black line) obtained from the run without AGN (top panel) or with AGN (middle panel).
+In both runs, the total best-fit (black solid line) is obtained by adding up the attenuated stellar template (blue dot-dashed), the nebular emission (green lines), and the AGN (red dashed).
+The AGN template shown in Fig.
+3 breaks down into an accretion disk component (yellow dot-dashed) peaking in the optical-NIR, and a dusty component (clumpy torus and polar dust; red dot-dashed) peaking at longer wavelengths.
+We further note that the best-fit stellar template chosen in the AGN run has very little dust attenuation (AV0.01) and the far-IR emission is therefore much fainter than the ALMA upper limit and entirely attributed to AGN (polar) dust.
+Regardless of the origin of Optical/UV emission in LRDs (see Sect.
+4.1.1 for more details), we argue that AGN-heated dust is needed to reproduce the monothonic increase in the rest 1–3m, which pure dust-attenuated stellar light fails to match.
+This is further supported by our ALMA upper limit, which is very similar to that derived by Casey et al.
+(2025) (see also Xiao et al.
+2025; Setton et al.
+2025), and that translates into Mdust106 M⊙ for T100 K.
+Indeed, the best-fit SED without AGN strongly overshoots the ALMA upper limit when attempting to reproduce the power-law-like SED at rest 1–3m.
+While this could be resolved by lifting the energy balance implemented in CIGALE, even ignoring the upper limits (at f2550w and ALMA/B6), a simple F-test comparing the two reduced values favours the AGN solution at 95% significance.
+4.1.1 Caveats and limitations of broad-band SED-fitting
+We caution the reader that our SED analysis does not endorse a pure stellar origin of the Optical/UV continuum, despite the stellar template chosen by CIGALE in the SED-fitting.
+We highlight some caveats and limitations inherent to our photometric SED analysis.
+For instance, CIGALE assumes a single-zone obscuration between optical, UV and IR light, which might be oversimplistic in case of no co-spatiality (see Sect.
+4.3).
+In addition, broad emission lines - as those observed in most LRD spectra - are not properly plugged-in the CIGALE library.
+And more importantly, despite our median stacking mitigates broad emission line contamination, it relies on broad-band photometry, which most likely irons out the strongest/steepest Balmer breaks, with potential impact on the output stellar ages and .
+Regarding limitations on the AGN templates, instead, while a “BH-Star” model (Naidu et al.
+2025) could be better suited for reproducing strong Balmer breaks and broad emission lines (e.g.
+de Graaff et al.
+2025), a satisfactory fit of the full SED (as shown in Fig.
+3, bottom panel) further requires a dusty torus and (in most cases) an additional UV component.
+In this framework, accommodating a joint BH-Star and dusty-torus template demands radiative transfer modeling (following e.g.
+skirtor; Stalevski et al.
+2012; 2016).
+Therefore, a more faithful modeling of the spectro-photometric LRD features would likely requires an ad-hoc grid of (AGN) models with those built-in features.
+Besides the uncertainty related to representativeness of such realistic models, this task is beyond the scope of this paper, which focuses on hot-dust emission from stacking photometric data.
+4.2 Comparison with literature LRD SEDs
 In the literature, LRDs are inevitably selected with distinct criteria, depending on the availability of multi-band NIRCam and MIRI imaging, and rest-frame optical spectra.
 While our LRD sample meets the same compactness and V-shape criteria across all fields, we compare our stacked SED against other LRD SEDs based on different selections.
-The fourth panel of Fig.
-5 shows for comparison: the median stacked SED from Akins et al.
-(2024; green symbols); the corresponding galaxy and AGN best-fit LRD SED from Casey et al.
-(2024; dot-dashed lines); the ALMA upper limit from Casey et al.
-(2025; downward triangle) and the observed CAPER-z9 NIRSpec spectrum (at =9.288; Taylor et al.
-2025) shifted to =6.2 (orange symbols).
+Fig.
+4 displays for comparison a number of stacked and single LRD SEDs, all in the rest-frame and normalized to the rest 1 m flux density to emphasize the difference in NIR slope.
+Among the median stacked SEDs, we show: Akins et al.
+(2024; red starred symbols); the best-fit galaxy and AGN SED fits from Casey et al.
+(2024; dot-dashed lines); the median stacked SED obtained from the LRD sample of Kokorev et al.
+(2024; blue squares).
 Akins et al.
-(2024) selected 434 LRDs in the COSMOS-Web area (0.5 deg2) that simultaneously meet a compactness (f444w) cut, a very red colour cut (f444w/f277w) and a bad brown drarf fit, without explicitly requiring a “V-shaped” SED due to limitation in NIRCam imaging.
+(2024) selected 434 LRDs in the COSMOS-Web area (0.5 deg2; Casey et al.
+2023) that simultaneously meet a compactness (f444w) cut, a very red colour cut (f444w/f277w) and a bad brown dwarf fit, without explicitly requiring a “V-shaped” SED due to limitation in NIRCam imaging.
 Their median SED agrees remarkably well with ours, though their statistics in MIRI is limited to 7.7m (over 1/3 of COSMOS-Web) and 18m (from the inner PRIMER-COSMOS footprint, 4% of the full sample), this latter leading to tentative detection (0.430.15 mJy).
 We show that their 18m flux is fully consistent with the rising MIR slope seen longward 10m.
-Their upper limits are here re-scaled to 3 (Herschel and SCUBA limits are too shallow to enter the y-scale).
-Despite the different LRD selection and survey field (25% of our LRDs are in Akins et al.
+Their upper limits are here re-scaled to 3 Despite the different LRD selection and survey field (25% of our LRDs are in Akins et al.
 2024), this agreement further suggests that the compactness and the conservative colour selection by Akins et al.
 (2024) broadly compensate the need for a V-shaped SED, leading to self-consistent LRD samples.
-In the same fourth panel of Fig.
-5, we compare our stacks against the median LRD SED by Casey et al.
-(2024), that is inferred from a joint LRD sample (675 LRDs) from Akins et al.
+Kokorev et al.
+(2024) selected 260 LRDs over the same JWST Legacy fields of this study, based on a compactness (in f444w) criterion and various colour cuts that mimic a V-shape requirement (see their Sect.3.1 in Kokorev et al.
+2024 for details).
+From their list, we further update the redshift measurements with additional as done in Sect.
+2, and repeat the same median stacking analysis carried out for our LRD sample, finding good consistency.
+In Fig.
+4 we also compare our stacks against the median best-fit SED by Casey et al.
+(2024), that was obtained from stacking a joint LRD sample (675 LRDs) from Akins et al.
 (2024) and Kokorev et al.
 (2024).
-This latter consists of a compactness (in f444w) criterion and various colour cuts that mimic a V-shape requirement (see their Sect.3.1 in Kokorev et al.
-2024 for details).
 In their work, Casey et al.
-(2024) performed stacking in rest-frame wavelength bins, following the observed SED of each LRD.
-The median stacked SED was then redshifted to =6.2 (as for our sample) and fitted with either a pure AGN (red) or a pure galaxy template (blue).
+(2024) performed stacking in rest-frame wavelength bins, applying a K-correction based on the observed SED of each LRD.
+The median stacked SED was then redshifted to =6.2 (as for our sample) and fitted with either a pure AGN (dashed red) or a pure galaxy template (dashed blue).
 We emphasize that our median stacks are fully consistent with their median AGN SED within their uncertainties (here not reported).
-If any, the AGN SED from Casey et al.
-(2024) stands slightly higher than our MIRI points, that partly reflects the MIRI upper limits in their study (20 objects per bin at 0.8m).
-Our stacks and their SED are also mutually consistent with the ALMA upper limit derived in Casey et al.
-(2025) by stacking 60 LRDs (among the Akins et al.
+The extrapolation of our AGN SED at mm wavelengths is also consistent (i.e.
+below) with the ALMA upper limit derived in Casey et al.
+(2025) from stacking 60 LRDs (among the Akins et al.
 2024 sample) at =1.3 mm.
-In addition, we compare our stacks against the NIRSpec spectrum of CAPER-z9 at =9.288 (Taylor et al.
-2025), which is complemented by MIRI data at 10m (detection) and 21m (upper limit).
-This continuous spectrum enables a direct comparison with our SED shape over a broad wavelength range.
-After moving the CAPER-z9 spectrum and photometry to =6.2 (orange symbols), we again find a reassuring consistency with our data.
-Another meaningful comparison with MIRI studies of LRDs is against the sample of Williams et al.
+In addition, we display for comparison the spectra and photometric SEDs of some well-studied LRDs.
+These comprise: CAPER-z9 (=9.288; Taylor et al.
+2025); A2744-45924 (=4.4655; Labbe et al.
+2024); RUBIES-BLAGN-1 (=3.1034; Wang et al.
+2025); the “Big Red Dot” (BiRD, =2.33; Loiacono et al.
+2025); the “Rosetta Stone” (or “GN 280754”; =2.26; Juodžbalis et al.
+2024a) and “The Cliff” (=3.55; de Graaff et al.
+2025).
+This compilation showcases the variety of rest-frame 1–3 m SEDs of bright LRDs, as compared to median stacked SEDs.
+As noted here and in other works (e.g.
+Barro et al.
+2024; Loiacono et al.
+2025; Setton et al.
+2025; Ronayne et al.
+2025), most of the brightest LRDs do not seem to show a rising NIR slope but a flat, or even decreasing, rest-frame 1–3 m SEDs.
+This is at odds with recent MIRI stacked SEDs, which are dominated by non-detections (only 10% are detected longward f1000w, see Appendix A.2).
+Such a comparison further highlights that LRDs are not a single homogeneous population, and that our median stacking in the longest MIRI bands is more representative of intrinsically fainter sources.
+Another meaningful comparison with MIRI studies of LRDs comprises the work of Williams et al.
 (2024), who stacked LRDs in all MIRI bands within the GOODS-SMILES footprint.
@@
 2025), albeit they are certainly an interesting sub-sample to follow-up.
-B.5 Effect of a broad redshift range
+4.3 Are LRDs typically dust-obscured AGN?
+By considering the best-fit solution with AGN, the rest-frame 6 m luminosity of the AGN torus can be used as a proxy for the intrinsic AGN X-ray (rest [2–10] keV) luminosity (X,intr (e.g.
+Stern 2015).
+From 6μm=(3.80.6)1043 erg/s, we derive X,intr=21043 erg/s, which nicely matches that expected from the best-fit AGN bolometric luminosity, if assuming standard X-ray bolometric corrections (e.g.
+Lusso et al.
+2012).
+By stacking the deepest X-ray images of the 44 LRDs falling inside the Chandra Deep Field South (Luo et al.
+2017), we obtain a 3 [2–10] keV luminosity limit X1042 erg/s at =6.2 (Comastri et al.
+2025).
+In order to match it with the X,intr expected from the AGN SED, the hydrogen column density has to be NH31024 cm-2, implying Compton-thick gas obscuration (Ananna et al.
+2024; Yue et al.
+2024; Maiolino et al.
+2025).
+Given the SED-based estimate of the AGN bolometric luminosity, erg s-1 and the best-fit stellar mass from the AGN solution (M⋆2.5109 M⊙; i.e.
+2 smaller than from a pure-galaxy fit), we emphasize that, if LRDs are highly-accreting AGN (Eddington ratio 10%), the BH-to-galaxy mass ratio MBH/M⋆1:100, implying strongly overmassive BHs compared to local scaling relations (e.g.
+Reines & Volonteri 2015).
+Moreover, our estimate corresponds to a dust sublimation radius of 0.1 pc at K (Barvainis 1987), meaning that any dusty AGN structure should begin at scales pc from the central SMBH.
+In the bottom panel of Fig.
+3, we try to decompose the AGN-dust template with three multi-temperature greybody curves (with free to vary).
+Our rising MIRI SED up to rest-frame 2–3m indicates hot dust at K, though it could be even hotter if optically-thick.
+Instead, the bulk dust mass (Mdust3104 M⊙) comes from colder polar AGN dust at 50 K (though not constrained by our data).
+While a dusty structure could induce X-ray obscuration, it would also strongly suppress the Optical/UV continuum, in tension with the blue UV slope and strong Balmer break features that imply relatively modest dust attenuation (; Akins et al.
+2024; Casey et al.
+2025).
+This apparent contradiction can be resolved if X-ray absorption in LRDs occurs within (0.1 pc), in a gas-dense and dust-free region, implying no co-spatiality.
+We stress that, despite our stacking analysis suggests that most (50%) LRDs contain hot-AGN dust, they do not need to be dust-obscured.
+This gas-dust displacement implies that theoretical models of Quasi-star (Begelman & Dexter 2025), BH–star (Naidu et al.
+2025), and super-massive star (Nandal & Loeb 2025), can in principle co-exist with outer hot dust, albeit its formation scenario (in-situ or ex-situ) remains open.
+Our findings also fit with an alternative hypothesis (Setton et al.
+2024; Maiolino et al.
+2025; D’Eugenio et al.
+2025a; Ji et al.
+2025; Inayoshi & Maiolino 2025), postulating that X-ray obscuration arises from Compton-thick gas clouds within the BLR (at sub, i.e.
+dust-free), which covers the compact X-ray corona without suppressing the broad Balmer lines.
+Yet, the origin of the Optical/UV continuum remains open.
+In this debated framework, recent works report the presence of a forest of auroral and FeII lines in some luminous LRDs (e.g.
+Tripodi et al.
+2025; Lin et al.
+2025; Lambrides et al.
+2025; Torralba et al.
+2025; D’Eugenio et al.
+2025b), which indicate that - at least in these bright objects - part of the UV continuum is dominated by a compact AGN embedded in a dense gas cocoon, that (at least partly) explains the Balmer break and the broadening of the Balmer lines.
+Regardless of the AGN orientation, we argue that the hot-dust revealed in this work should be placed most likely at outer radii (Rsub0.1 pc) than the BLR scales at which the (dust-free) gas absorption takes place.
+Given that Rsub, we expect that even in the most luminous LRDs (e.g.
+Torralba et al.
+2025; D’Eugenio et al.
+2025b, i.e.
+50-100 brighter than our median stacked AGN), Rsub1 pc, that is 104 times larger than the nominal thickness of the BLR (gas) cloud (D’Eugenio et al.
+2025b).
+Although we caution that LRDs might not consist of a single population, we speculate two simple scenarios in Fig.
+5, depending on what powers the Optical/UV continuum in LRDs.
+In either cases, the observed Optical/UV continuum comes from dust-unattenuated regions, despite the presence of hot AGN dust, and the outer BLR is directly visible along the line of sight.
+(i) If the Optical/UV comes from unobscured stellar light (Fig.
+5, left), it must be placed at larger scales than the AGN torus.
+In this case the LRD resembles a “Type-2” AGN configuration, in which the torus covers the AGN accretion disk, which would otherwise outshine the host.
+In case of high torus clumpiness, mildly-attenuated light from the outer accretion disk and the BLR could pierce through the torus and add to the optical/UV emission.
+(ii) Instead, if the Optical/UV is dominated by AGN light (Fig.
+5, right), the LRD resembles a “Type-1” unobscured AGN configuration, though the origin of the sharp Balmer break remains unclear.
+Fainter Optical/UV light arising from (kpc-scale) star formation cannot be ruled out (e.g.
+Rinaldi et al.
+2025).
+5 Conclusions
+In this manuscript, we investigate the nature of LRDs via an extensive NIRCam, MIRI (and ALMA) stacking analysis of a homogeneously-selected sample of 302 sources across the CEERS, JADES, and PRIMER JWST fields.
+We fit the median stacked SED with either pure galaxy or galaxy+AGN templates to interpret the average IR emission properties of LRDs.
+In summary, we find clear evidence for a rising MIRI continuum up to 3m, which we interpret via SED fitting as hot (820 K) AGN dust, regardless of whether the Optical/UV continuum is stellar- or AGN-dominated, or a combination thereof.
+Despite the presence of hot dust, these LRDs selected via V-shaped and compactness criteria must be not strongly dust-obscured.
+The non-detection from deep X-ray stacking suggest that the apparent X-ray weakness of LRDs is caused by Compton-thick (31024 cm-2) gas obscuration.
+We acknowledge that our homogeneous LRD selection does not necessarily translate into a single LRD population with similar physical properties.
+However, we reiterate that our main empirical finding is unchanged, that is, the majority () of LRDs show AGN-heated dust emission.
+If this were not the case, the rising NIR slope would not appear in the median stacked data, and the mean and median MIRI fluxes would differ significantly after removing individual detections (contrary to our findings in Appendix A.2).
+The small fraction of individual MIRI detections (10% longward f1000w, i.e.
+longer than rest 1.3m) contributes negligibly to the median stacked SED at rest-frame 1–3m.
+Therefore, the rising NIR slope is predominantly driven by the MIRI non-detections, suggesting that AGN-heated dust is characteristic of the fainter, more typical LRD population.
+That being said, our results are contingent upon the adopted LRD selection criteria, which may be inherently biased.
+For instance, if the rest-frame UV emission of LRDs arises from an AGN accretion disk, we speculate that photometric LRD selection criteria (especially the blue UV-slope requirement) would preferentially select “Type-1” AGN, according to the Unified Model.
+The same criterion would therefore induce a bias against the “Type-2” LRD analogs, in which the disk is hidden by an obscuring dusty structure.
+With this caveat in mind, our results should be taken as an empirical benchmark for motivating deeper investigations of the full LRD population.
+Acknowledgements.
+The authors thank the anonymous referee for their constructive report.
+ID thanks Federica Loiacono for sharing their SED datapoints, and also Yoshiki Matsuoka and Guillermo Barro for insightful discussions.
+ID acknowledges funding by the European Union – NextGenerationEU, RRF M4C2 1.1, Project 2022JZJBHM: ”AGN-sCAN: zooming-in on the AGN-galaxy connection since the cosmic noon” - CUP C53D23001120006.
+(Some of) the data producs presented here were retrieved from the DJA, an initiative of the Cosmic Dawn Center (DAWN), which is funded by the Danish National Research Foundation under grant DNRF140.
+We acknowledge the contribution of the COSMOS collaboration, consisting of more than 200 scientists.
+More information about the COSMOS survey can be found at https://cosmos.astro.caltech.edu/.
+This work is based [in part] on observations made with the NASA/ESA/CSA James Webb Space Telescope.
+The data were obtained from the Mikulski Archive for Space Telescopes at the Space Telescope Science Institute, which is operated by the Association of Universities for Research in Astronomy, Inc., under NASA contract NAS 5-03127 for JWST.
+These observations are associated with program 1180, 1181, 1210, 1286, 1895, 1963, 3215, 1345, 1837 and 1207.
+The data described here may be obtained from the MAST archive at https://doi.org/10.17909/et3f-zd57 (SMILES) and https://dx.doi.org/10.17909/8tdj-8n28 (JADES) and on the DJA at https://s3.amazonaws.com/grizli-v2/JwstMosaics/v7/index.html (CEERS and PRIMER-UDS).References
+- Adscheid et al. (2024) Adscheid, S., Magnelli, B., Liu, D., et al. 2024, A&A, 685, A1
+- Akins et al. (2024) Akins, H. B., Casey, C. M., Lambrides, E., et al. 2024, arXiv e-prints, arXiv:2406.10341
+- Alberts et al. (2024) Alberts, S., Lyu, J., Shivaei, I., et al. 2024, ApJ, 976, 224
+- Ananna et al. (2024) Ananna, T. T., Bogdán, Á., Kovács, O. E., Natarajan, P., & Hickox, R. C. 2024, ApJ, 969, L18
+- Barro et al. (2024) Barro, G., Perez-Gonzalez, P. G., Kocevski, D. D., et al. 2024, arXiv e-prints, arXiv:2412.01887
+- Barvainis (1987) Barvainis, R. 1987, ApJ, 320, 537
+- Begelman & Dexter (2025) Begelman, M. C. & Dexter, J. 2025, arXiv e-prints, arXiv:2507.09085
+- Boquien et al. (2019) Boquien, M., Burgarella, D., Roehlly, Y., et al. 2019, A&A, 622, A103
+- Brammer (2023a) Brammer, G. 2023a
+- Brammer (2023b) Brammer, G. 2023b
+- Bruzual & Charlot (2003) Bruzual, G. & Charlot, S. 2003, MNRAS, 344, 1000
+- Carranza-Escudero et al. (2025) Carranza-Escudero, M., Conselice, C. J., Adams, N., et al. 2025, arXiv e-prints, arXiv:2506.04004
+- Casey et al. (2025) Casey, C. M., Akins, H. B., Finkelstein, S. L., et al. 2025, arXiv e-prints, arXiv:2505.18873
+- Casey et al. (2024) Casey, C. M., Akins, H. B., Kokorev, V., et al. 2024, ApJ, 975, L4
+- Casey et al. (2023) Casey, C. M., Kartaltepe, J. S., Drakos, N. E., et al. 2023, ApJ, 954, 31
+- Cenci & Habouzit (2025) Cenci, E. & Habouzit, M. 2025, MNRAS, 542, 2597
+- Chabrier (2003) Chabrier, G. 2003, PASP, 115, 763
+- Charlot & Fall (2000) Charlot, S. & Fall, S. M. 2000, ApJ, 539, 718
+- Chen et al. (2025) Chen, K., Li, Z., Inayoshi, K., & Ho, L. C. 2025, arXiv e-prints, arXiv:2505.22600
+- Comastri et al. (2025) Comastri, A., Lanzuisi, G., Vito, F., et al. 2025, arXiv e-prints, arXiv:2510.00112
+- Dale et al. (2014) Dale, D. A., Helou, G., Magdis, G. E., et al. 2014, ApJ, 784, 83
+- de Graaff et al. (2024) de Graaff, A., Rix, H.-W., Carniani, S., et al. 2024, A&A, 684, A87
+- de Graaff et al. (2025) de Graaff, A., Rix, H.-W., Naidu, R. P., et al. 2025, arXiv e-prints, arXiv:2503.16600
+- D’Eugenio et al. (2025a) D’Eugenio, F., Nelson, E., Ji, X., et al. 2025a, arXiv e-prints, arXiv:2510.00101
+- D’Eugenio et al. (2025b) D’Eugenio, F., Nelson, E., Ji, X., et al. 2025b, arXiv e-prints, arXiv:2510.00101
+- Eisenstein et al. (2023) Eisenstein, D. J., Johnson, B. D., Robertson, B., et al. 2023, arXiv e-prints, arXiv:2310.12340
+- Finkelstein et al. (2023) Finkelstein, S. L., Bagley, M. B., Ferguson, H. C., et al. 2023, ApJ, 946, L13
+- Franco et al. (2025) Franco, M., Casey, C. M., Koekemoer, A. M., et al. 2025, arXiv e-prints, arXiv:2506.03256
+- Gloudemans et al. (2025) Gloudemans, A. J., Duncan, K. J., Eilers, A.-C., et al. 2025, ApJ, 986, 130
+- Greene et al. (2024) Greene, J. E., Labbe, I., Goulding, A. D., et al. 2024, ApJ, 964, 39
+- Greene et al. (2025) Greene, J. E., Setton, D. J., Furtak, L. J., et al. 2025, arXiv e-prints, arXiv:2509.05434
+- Hainline et al. (2025) Hainline, K. N., Maiolino, R., Juodžbalis, I., et al. 2025, ApJ, 979, 138
+- Harish et al. (2025) Harish, S., Kartaltepe, J. S., Liu, D., et al. 2025, arXiv e-prints, arXiv:2506.03306
+- Heintz et al. (2025) Heintz, K. E., Brammer, G. B., Watson, D., et al. 2025, A&A, 693, A60
+- Hviding et al. (2025) Hviding, R. E., de Graaff, A., Miller, T. B., et al. 2025, arXiv e-prints, arXiv:2506.05459
+- Inayoshi & Maiolino (2025) Inayoshi, K. & Maiolino, R. 2025, ApJ, 980, L27
+- Inayoshi et al. (2025) Inayoshi, K., Murase, K., & Kashiyama, K. 2025, arXiv e-prints, arXiv:2509.19422
+- Inayoshi et al. (2020) Inayoshi, K., Visbal, E., & Haiman, Z. 2020, ARA&A, 58, 27
+- Ji et al. (2025) Ji, X., Maiolino, R., Übler, H., et al. 2025, arXiv e-prints, arXiv:2501.13082
+- Juodžbalis et al. (2024a) Juodžbalis, I., Ji, X., Maiolino, R., et al. 2024a, MNRAS, 535, 853
+- Juodžbalis et al. (2024b) Juodžbalis, I., Maiolino, R., Baker, W. M., et al. 2024b, Nature, 636, 594
+- Juodžbalis et al. (2025) Juodžbalis, I., Marconcini, C., D’Eugenio, F., et al. 2025, arXiv e-prints, arXiv:2508.21748
+- King (2025) King, A. 2025, MNRAS, 536, L1
+- Kocevski et al. (2025) Kocevski, D. D., Finkelstein, S. L., Barro, G., et al. 2025, ApJ, 986, 126
+- Kokorev et al. (2024) Kokorev, V., Caputi, K. I., Greene, J. E., et al. 2024, ApJ, 968, 38
+- Labbe et al. (2025) Labbe, I., Greene, J. E., Bezanson, R., et al. 2025, ApJ, 978, 92
+- Labbe et al. (2024) Labbe, I., Greene, J. E., Matthee, J., et al. 2024, arXiv e-prints, arXiv:2412.04557
+- Labbé et al. (2023) Labbé, I., van Dokkum, P., Nelson, E., et al. 2023, Nature, 616, 266
+- Lambrides et al. (2024) Lambrides, E., Garofali, K., Larson, R., et al. 2024, arXiv e-prints, arXiv:2409.13047
+- Lambrides et al. (2025) Lambrides, E., Larson, R., Hutchison, T., et al. 2025, arXiv e-prints, arXiv:2509.09607
+- Leung et al. (2024) Leung, G. C. K., Finkelstein, S. L., Pérez-González, P. G., et al. 2024, arXiv e-prints, arXiv:2411.12005
+- Lin et al. (2025) Lin, X., Fan, X., Wang, F., et al. 2025, arXiv e-prints, arXiv:2504.08039
+- Liu et al. (2019) Liu, D., Schinnerer, E., Groves, B., et al. 2019, ApJ, 887, 235
+- Liu et al. (2025) Liu, H., Jiang, Y.-F., Quataert, E., Greene, J. E., & Ma, Y. 2025, arXiv e-prints, arXiv:2507.07190
+- Loiacono et al. (2025) Loiacono, F., Gilli, R., Mignoli, M., et al. 2025, arXiv e-prints, arXiv:2506.12141
+- Luo et al. (2017) Luo, B., Brandt, W. N., Xue, Y. Q., et al. 2017, ApJS, 228, 2
+- Lusso et al. (2012) Lusso, E., Comastri, A., Simmons, B. D., et al. 2012, MNRAS, 425, 623
+- Madau (2025) Madau, P. 2025, arXiv e-prints, arXiv:2501.09854
+- Madau & Haardt (2024) Madau, P. & Haardt, F. 2024, ApJ, 976, L24
+- Magnelli et al. (2024) Magnelli, B., Adscheid, S., Wang, T.-M., et al. 2024, A&A, 688, A55
+- Maiolino et al. (2025) Maiolino, R., Risaliti, G., Signorini, M., et al. 2025, MNRAS, 538, 1921
+- Maiolino et al. (2024) Maiolino, R., Scholtz, J., Curtis-Lake, E., et al. 2024, A&A, 691, A145
+- Matthee et al. (2024) Matthee, J., Naidu, R. P., Brammer, G., et al. 2024, ApJ, 963, 129
+- Mazzolari et al. (2025) Mazzolari, G., Scholtz, J., Maiolino, R., et al. 2025, A&A, 700, A12
+- Naidu et al. (2025) Naidu, R. P., Matthee, J., Katz, H., et al. 2025, arXiv e-prints, arXiv:2503.16596
+- Nandal & Loeb (2025) Nandal, D. & Loeb, A. 2025, arXiv e-prints, arXiv:2507.12618
+- Pacucci & Narayan (2024) Pacucci, F. & Narayan, R. 2024, ApJ, 976, 96
+- Pacucci et al. (2023) Pacucci, F., Nguyen, B., Carniani, S., Maiolino, R., & Fan, X. 2023, ApJ, 957, L3
+- Pérez-González et al. (2024) Pérez-González, P. G., Barro, G., Rieke, G. H., et al. 2024, ApJ, 968, 4
+- Reines & Volonteri (2015) Reines, A. E. & Volonteri, M. 2015, ApJ, 813, 82
+- Rieke et al. (2024) Rieke, G. H., Alberts, S., Shivaei, I., et al. 2024, ApJ, 975, 83
+- Rinaldi et al. (2025) Rinaldi, P., Rieke, G. H., Wu, Z., et al. 2025, arXiv e-prints, arXiv:2507.17738
+- Ronayne et al. (2025) Ronayne, K., Papovich, C., Kirkpatrick, A., et al. 2025, arXiv e-prints, arXiv:2508.20177
+- Rusakov et al. (2025) Rusakov, V., Watson, D., Nikopoulos, G. P., et al. 2025, arXiv e-prints, arXiv:2503.16595
+- Sacchi & Bogdán (2025) Sacchi, A. & Bogdán, Á. 2025, ApJ, 989, L30
+- Setton et al. (2024) Setton, D. J., Greene, J. E., de Graaff, A., et al. 2024, arXiv e-prints, arXiv:2411.03424
+- Setton et al. (2025) Setton, D. J., Greene, J. E., Spilker, J. S., et al. 2025, arXiv e-prints, arXiv:2503.02059
+- Stalevski et al. (2012) Stalevski, M., Fritz, J., Baes, M., Nakos, T., & Popović, L. Č. 2012, MNRAS, 420, 2756
+- Stalevski et al. (2016) Stalevski, M., Ricci, C., Ueda, Y., et al. 2016, MNRAS, 458, 2288
+- Stern (2015) Stern, D. 2015, ApJ, 807, 129
+- Taylor et al. (2025) Taylor, A. J., Kokorev, V., Kocevski, D. D., et al. 2025, ApJ, 989, L7
+- Torralba et al. (2025) Torralba, A., Matthee, J., Pezzulli, G., et al. 2025, arXiv e-prints, arXiv:2510.00103
+- Tripodi et al. (2025) Tripodi, R., Bradač, M., D’Eugenio, F., et al. 2025, arXiv e-prints, arXiv:2507.20684
+- Valentino et al. (2023) Valentino, F., Brammer, G., Gould, K. M. L., et al. 2023, ApJ, 947, 20
+- Valentino et al. (2025) Valentino, F., Heintz, K. E., Brammer, G., et al. 2025, A&A, 699, A358
+- Volonteri et al. (2021) Volonteri, M., Habouzit, M., & Colpi, M. 2021, Nature Reviews Physics, 3, 732
+- Wang et al. (2025) Wang, B., de Graaff, A., Davies, R. L., et al. 2025, ApJ, 984, 121
+- Wang et al. (2024) Wang, T.-M., Magnelli, B., Schinnerer, E., et al. 2024, A&A, 681, A110
+- Wang et al. (2022) Wang, T.-M., Magnelli, B., Schinnerer, E., et al. 2022, A&A, 660, A142
+- Williams et al. (2024) Williams, C. C., Alberts, S., Ji, Z., et al. 2024, ApJ, 968, 34
+- Xiao et al. (2025) Xiao, M., Oesch, P. A., Bing, L., et al. 2025, arXiv e-prints, arXiv:2503.01945
+- Yue et al. (2024) Yue, M., Eilers, A.-C., Ananna, T. T., et al. 2024, ApJ, 974, L26
+Appendix A Stacking tests
+We test our stacking results against multiple potential biases related to sample selection, stacking method, field-to-field variations and photometric redshift uncertainties.
+The main outcome of these tests is shown in Fig.
+6.
+In general, these tests further corroborate the robustness of our stacking results and the need for an obscured AGN component to reproduce the MIR excess.
+Each of the following tests is discussed in a dedicated subsection, numbered as the panels of Fig.
+6 from top to bottom.
+A.1 Mean vs median stacking
+In the first panel of Fig.
+6 we compare our median stacked fluxes (grey open circles) against mean stacked fluxes (red starred symbols).
+These latter values are computed from the inverse-variance weighted flux density map, accounting for the different sensitivity (i.e.
+exposure time) of each target.
+Mean stacking leads to higher fluxes with respect to median stacking.
+This is likely caused by the fact that the resulting (rms-weighted) linear mean flux can be boosted upwards by relatively bright individual detections (we note that upper limits at f2550w and ALMA/B6 are unchanged as they are calculated from the rms map).
+This is the case in NIRCam bands, where all LRDs are individually detected by definition over multiple bands, both blueward and redward the Balmer break (in order to assess the V-shape).
+An even stronger boost is seen at MIRI 7.7m and 18m, which count about 100 more targets than at other MIRI bands (see Table 1 and Appendix A.2).
+In these two MIRI bands, we investigate the origin of these fluctuations, finding that the f770w and f1800w fluxes are over-boosted relative to other bands because of two strong outliers at 4.5 in the PRIMER-UDS field.
+These are: RUBIES-BLAGN1 (or PRIMER-UDS 38223, at =spec=3.1; Wang et al.
+2025), which is a bright X-ray detected BLAGN; and PRIMER-UDS 9235 (=phot=4.45), previously flagged in Leung et al.
+(2024) to be the strongest f1800w emitter of the PRIMER-UDS field.
+We check that after removing only those two PRIMER-UDS outliers, the ratio between mean and median stacked fluxes (over all fields) re-aligns to the average factor of 2-3 seen in all other bands.
+We do not remove these outliers in the rest of our work, in order to not fine-tune the results.
+However, this is again a caveat on using mean stacks without pre-removing bright individual detections.
+Moreover, we opt for median stacking also because it strongly mitigates contamination from bright neighbors, and thus reduces the confusion noise for the faint sources.
+This comparison, however, demonstrates that mean and median stacked SEDs display broadly-similar shapes.
+A.2 Contribution of MIRI detections to the stacking
+We test the robustness of our stacking results against the contribution of individual detections.
+We repeat both mean (open stars) and median (open circles) stacking after removing individual MIRI detections (with S/N3), in each band.
+In Figure 6 (2nd panel from top) we display the new stacking results.
+The individual MIRI detections are shown as small blue circles, and the corresponding detection rate is reported at each MIRI band.
+We notice that mean and median stacking of MIRI non-detections lead to well-consistent results, also in f560w and f770w where the detection rate is above 70%.
+In addition, for MIRI non-detections we still observe a significant rise at rest-frame 1–3 m, while individual MIRI detections stand out at much higher fluxes.
+This check strongly suggests that the difference between median and mean stacks boils down to the number of individual detections, and that the rising near-IR slope is driven by the bulk of non-detections, which is sensitive to more common and intrinsically fainter sources.
+A.3 Field-to-field variations
+An important aspect that our analysis relies on is the combination of multiple NIRCam and MIRI survey fields.
+While this is necessary to improve the statistics and the stacked S/N, it might potentially introduce systematics related to different JWST coverage and depth, which could affect the properties of photometrically-selected LRDs.
+We thus verify that our stacking results are robust against field-to-field variations, by repeating the same analysis separately for each field.
+The third panel in Fig.
+6 displays the median stacked SED obtained in each field (represented by different filled symbols) overlaid to our full stacking results (grey open circles).
+For visual purposes, here we relax the detection threshold to S/N2 (hence upper limit fluxes are set to twice the noise level if S/N2).
+Despite their heterogeneous spectral coverage across MIRI bands, all single-field median SEDs are consistent with the combined median SED.
+Moreover, we verify that stacking the same exact 22 objects having simultaneously NIRCam and MIRI coverage in all bands (from the GOODS-S field), the resulting SED is fully consistent with that obtained via stacking the full LRD sample.
+These checks strengthen that similarly-selected LRDs across different fields share a common median SED, despite their and heterogeneous NIRCam/MIRI spectral coverage.
+A.4 Effect of photometric redshifts
+A relevant caveat in our analysis is the photometric selection of LRDs.
+Specifically, photometric redshift uncertainties can be notably high, potentially biasing the LRD selection.
+We iterate that for the spectroscopic subsample (96/302 objects) there is good agreement between photometric and spectroscopic redshifts (Sect.
+2).
+However, we test this directly by stacking NIRCam and MIRI images solely for the spectroscopic subsample.
+The fourth panel in Fig.
+6 displays the corresponding median stacked SED (blue filled circles).
+Numbers are given in Table 1 for each field: when combining all fields there are between 36 and 53 objects in MIRI (except for f560w, f1280w and f2550w with 5 objects due to a smaller areal coverage) and 96 objects in NIRCam.
+Despite the S/N being smaller than that of the full sample, the two median stacked SEDs agree remarkably well with each other.
+We also explore whether a cleaner selection of “bona fide” AGN based on broad emission lines yields similar trends.
+From the NIRSpec Merged Table of the DJA, we isolate 57/96 (59%) objects with broad Hα or Hβ emission lines, as inferred from the best-fit continuum fluxes, line intensity and equivalent widths.
+This fraction agrees with that reported in the literature (e.g.
+Greene et al.
+2024; Hviding et al.
+2025).
+The corresponding median stacked SED of BL AGN (empty starred symbols), displays no significant difference with respect to the full spec- (96) subsample, albeit a lower S/N in all bands due to smaller statistics.
+A.5 Effect of a broad redshift range
 As shown in Fig.
-4, our targets span a very broad redshift range, from 2 to 11, with over 90% of the sample being within the range (48).
+1, our targets span a very broad redshift range, from 2 to 11, with over 90% of the sample being within the range (48).
 Because we perform stacking in the observed frame, the contribution to each stacked point comes from very different parts of the rest-frame spectrum.
@@
 We show this comparison in the bottom panel of Fig.
-5.
+6.
 The lower- bin shows slightly brighter fluxes and steeper MIRI slope compared to the higher- bin, as expected given the smaller distance and the longer rest-frame wavelength probed (e.g.
 3.5 m instead of 2.5 m at MIRI-21 m), respectively.
+We also note that fluxes are in energy units, hence the apparent flattening of the 68 NIR SED translates into a monothonic rise in flux density () units, as seen in the full sample.
 Overall, this check suggests that LRDs selected at different redshifts have broadly-consistent stacked SEDs, with no evidence for a galaxy-like drop at rest 1–2 m.
@@
 Given the broad similarity of the two SEDs in two different -bins, we believe this is unlikely the case.
-Appendix C Input SED-fitting parameters
+Appendix B Input SED-fitting parameters
 In the following Table, we outline the main input parameters used for SED-fitting with CIGALE (Section 4.1).
 | Module | Parameter | Symbol | Values |
-[…]
+| Stellar -folding time | ( yr) | 150, 200, 300, 500, 700, 900∗ | |
 | Star formation history | Stellar age | ( yr) | 200, 300, 500, 700, 900∗ |
@@
 | SFR ratio after/before the burst | 0.1, 0.5, 1∗, 2, 5, 10 ,30, 50 | ||
-[…]
+| Single stellar population | Initial mass function | – | Chabrier (2003) |
 | (Bruzual & Charlot 2003) | Metallicity | 0.0004, 0.008, 0.02∗ | |
-[…]
+| Nebular emission | Ionization parameter | -2.0∗, -1.0 | |
 | Gas metallicity | Zgas | 0.0004∗, 0.008, 0.019 | |
 | Electron density | 100,1000∗ | ||
-[…]
+| V-band attenuation (ISM) | AV,ISM | 0.01∗, 0.1, 0.3, 0.5, 0.7, 1.0, 1.5, 2, 2.5, 3, 3.5 | |
 | Dust attenuation | Fractional ISM attenuation | AV,ISM / (AV,BC+AV,ISM) | 0.44∗ |
@@
 | Power-law slope (BC) | – | -1.3∗ | |
-[…]
+| Mass fraction of PAH | qpah | 0.47, 2.5, 3.9∗ | |
 | Galactic dust emission | Minimum radiation field | Umin | 1.0, 10, 20, 40∗ |
@@
 | Illuminated fraction [Umin–Umax] | 0.02∗ | ||
-[…]
+| Edge-on optical depth at 9.7m | 3∗, 7 | ||
 | Slope of radial dust density profile | – | 1.0∗ | |
@@
 | Emissivity of polar dust | – | 1.6∗ |
-Appendix D Output SED-fitting parameters
+Appendix C Output SED-fitting parameters
 | Parameter | Value and 1 error |